A Bibliography of Publications in Theoretical Computer Science: 2025–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

06 September 2025 Version 1.20

Title word cross-reference

(1 + ε)Δ [151]. (Δ + 1) [340]. (min, +) [171]. (n, f) [184]. (n, k) [192]. (P_{5}, H) [319]. 0, 1, 2 [47]. 1 [173]. 1/2 [300]. 2 [82, 3, 2, 337, 286, 219]. 3 [87, 86]. b [38]. D [341, 38]. Δ [340, 298]. G [83]. H [130]. K [330, 47, 193, 188, 63, 72, 88, 212, 331, 321]. L [296]. H [335]. O(n log_{2} n) [148]. P_{5} [139]. r [89]. t [289].

-approximation [300]. -ary [47]. -balanced [330]. -base [341]. -category [87]. -center [72]. -coloring [340]. -cyclic [219]. -edge-coloring [151]. -Edge-Connected [337, 286, 3, 2]. -error [47]. -evacuation [184]. -free [139, 130, 319]. -good-neighbor [83]. -hyperbolic [298]. -majority [86].

-median [331]. -packing [335]. -parity [212]. -planar [173]. -pseudoforest [89]. -relation [341]. -shortest [188]. -spanners [289]. -spin [82]. -star [192]. -supplier [321]. -synchronisation [212]. -systems [296]. -way [63]. -weight [88].

1 [249, 23, 56]. 1-extendability [351]. 11 [262, 177]. 12 [71, 45]. 13 [114, 292]. 17 [306]. 19 [81, 227, 344].

2 [58, 102]. 2023 [197]. 2025 [145, 153, 159, 249, 262, 272, 58, 71, 81, 92, 23, 45, 196, 205, 216, 227, 236, 102, 114, 125, 136, 168, 177, 186, 328, 344, 360, 284, 292, 306, 315]. 21 [125]. 22 [153, 205, 186]. 27 [315]. 28 [272, 92]. 29 [159, 136, 360]. 2d [65].

3 [168, 284]. 3/4 [350]. 30 [236].

4 [196].

7 [145, 216].

9 [328]. 943 [3, 286, 41]. 963 [337].

ABCD [68]. absolute [110]. accident [10]. action [55]. active [143]. acyclic [11]. Adaptive [62, 25, 112]. Adding [56]. Additional [210]. Additive [298]. addressing [104]. adversarial [324, 191]. AFA [255]. agent [75, 9, 253]. agents [312, 329]. agreement [50]. album [348]. Algebraic [40, 207]. algebras [217]. Algorithm [337, 286, 1, 41, 350, 316, 333, 298, 211, 3, 2, 256, 203, 208, 99, 214, 163, 89, 264, 321, 300, 10]. algorithmic [221]. Algorithms [240, 11, 285, 93, 265, 193, 334, 229, 64, 356, 27, 320, 34, 170, 288, 259, 63, 230, 330, 72, 95, 32, 246, 39, 266, 331, 78, 12, 98]. AlgoWin [197]. allocation [351, 143, 253]. allocations [66, 67, 251, 224]. almost [342]. alphabet [254, 342]. Alternating [241]. analog [55]. analysis [107, 346, 19]. analyze [238]. anonymity [141]. anonymous [54, 312, 25]. anti [244]. anti-Ramsey [244]. any [50]. applications [294]. approach [143, 67, 104, 191]. approaches [115]. approval [313]. approximate [116, 88]. Approximating [52, 8, 4]. Approximation [347, 337, 286, 27, 230, 331, 264, 90, 12, 98, 59, 350, 298, 3, 2, 64, 79, 88, 99, 321, 300, 336, 234]. April [145, 153, 159]. arbitrary [181]. arc [119]. arc-geodetic [119]. area [76, 181]. area-proportional [76]. areas [252]. argument [182]. arguments [179]. arithmetic [55, 308]. Arm [206]. arrangement [36]. arrays [311]. articulated [356]. Arto [187]. ary [47]. assisted [353, 7]. associated [198]. asymmetric [353]. Asynchronous [117, 154, 46, 245]. asynchrony [189, 34]. attractors [260]. attribute [7]. attribute-based [7]. auctions [143]. August [249, 262, 272]. autoencoder [318]. automata [166, 355, 210, 212, 104, 199]. automata-based [104]. automaticity [357]. automaton [342]. autonomous [285, 316, 165]. Availability [1, 41]. avoid [142]. Axiomatising [276]. axis [50].

B [282]. back [88]. balanced [330]. Balancing [61]. bandit [324]. base [341, 38]. Based [1, 41, 317, 117, 333, 303, 62, 211, 252, 42, 299, 67, 304, 231, 104, 190, 155, 97, 160, 7]. Bayesian [180]. before [213]. behaviors [104]. Best [206]. betting [174]. between [289]. beyond [133]. biclique [330, 88]. biclique-free [88]. Big [39]. bijective [73]. Bike [353]. Bike-assisted [353]. bimatrix [116]. binary [325, 290, 342, 214, 224]. bipartite [120, 338]. bisimulation [276]. bisubmodular [264]. black [18, 77]. block [303]. block-based [303]. blockchains [323]. blocking [10]. blue [137]. Blum [179]. Board [22, 44, 305, 314, 327, 343, 359, 57, 70, 80, 91, 101, 113, 124, 135, 144, 152, 158, 167, 176, 185, 195, 204, 215, 226, 235, 248, 261, 271, 283, 291]. Boolean [87]. bootstrapping [309]. boundaries [296]. bounded [82, 173, 188, 157, 299, 200, 267, 72, 338, 95, 280]. bounds [207, 147]. box [77]. branching [207]. Bribery [163]. Broadcast [257]. bubble [19]. bubble-sort [19]. budget [269]. budgeting [301]. Burn [293]. Burrows [73]. Byzantine [184, 189].

C [202]. C-shaped [202]. candidate [33]. capacitated [79, 234]. cardinality [51]. case [269]. catastrophic [323]. categorical [127]. category [87]. cats [310]. CCS [276]. cellular [210, 304]. center [178, 72]. centers [298]. central [53]. certification [139, 345]. chains [5]. chance [246]. chance-constrained [246]. change [250].

Channel [351]. channels [222]. characteristics [256]. Characterization [278]. characterizations [356]. charging [18, 239]. checking [299, 259, 183]. Chemical [55]. chemistry [127]. child [290]. chordal [6]. choreographic [228]. chores [67, 251, 224]. churn [323]. ciphertext [74]. circle [184, 54, 12]. circuits [87]. cl [201]. cl-Turing [201]. class [326]. classes [121, 232, 15, 335]. classification [39]. classifications [128]. claw [133]. claw-free [133]. cliques [170]. Closure [166, 341]. Cloud [1, 41, 303, 123, 7]. cloud-assisted [7]. cluster [117]. cluster-based [117]. clustered [99]. Clustering [336, 79, 72]. clusters [280]. CNN [318]. CNN-transformer [318]. co [220]. co-gem-free [220]. code [290]. codes [115]. cognitive [317]. collapse [112]. collapse-free [112]. collective [25]. collector [348]. color [274]. coloring [120, 232, 295, 151, 340, 175, 6]. colorings [170]. colourings [352]. Combinatorial [356, 143]. common [269, 288, 242]. communication [276, 353]. communities [68]. commuting [142]. Compact [179, 155]. Comparison [39, 83]. Compatibility [285]. Competitive [233]. complementation [129]. Complete [172, 310, 346, 49, 109, 98]. complete-transposition [346]. completeness [303]. completes [161]. completion [326]. Complexity [244, 69, 240, 82, 335, 148, 103, 313, 322, 345, 108, 66]. component [104]. component-based [104]. Compositional [237]. compressed [74]. compressive [318]. Comput [3]. computability [223]. Computable [128]. Computational [165, 108]. computations [55]. Computer [187, 286, 337, 41]. computers [55]. Computing [341, 130, 244, 46, 112]. concepts [317, 100]. concerning [26, 69]. concurrent [241]. conditional [36, 313, 111]. conditionally [277]. Conflict [120, 121]. Conflict-free [120, 121]. congruences [276]. Conjecture [56]. Connected [337, 286, 3, 2, 162, 307]. connectivity [36, 96, 111]. conquer [146]. consensus [189]. constrained [258, 246, 321]. constraint [51, 345, 336]. constraints [56]. constructed [182]. Constructing [18]. construction [202]. consumption [103]. contact [76]. containing [12]. containment [60]. contamination [324]. contextual [26]. contiguous [64]. continuous [128]. control [149, 313]. convergence [285]. conversions [142]. Convex [199, 280, 90, 270]. convex-hull [280]. convexity [240]. cop [173]. cop-number [173]. core [247]. Corrigendum [3, 337]. Cost [190]. Cost-sharing [190]. counting [31]. cover [93, 225, 4, 302, 137, 99, 300]. coverage [252]. covering [27, 339]. covers [130, 209]. Critical [319, 220]. crossing [334]. crowdsourcing [358]. cubes [49, 109]. curves [296]. cut [29, 338]. Cuts [310]. cycle [99, 98]. Cycles [260]. Cyclic [250, 349, 219]. cyclotomic [309].

DAG [107, 103]. data

[123, 61, 178, 32, 110, 39]. deadlocks [104]. Debt [84]. decentralized [117]. decidability [332]. decidable [342]. Deciding [219]. Decision [279, 166, 26, 214, 8, 160]. declined [19]. decomposable [171]. decomposition [172]. decompositions [94]. Deep [304]. defense [233]. definite [277]. degree [82, 157]. degrees [275]. Delaunay [169]. delays [63]. deletion [208, 89]. deletions [26]. delimiting [294]. delivery [9, 253]. dense [151, 275]. Density [47]. descriptional [108]. design [264, 358]. determination [313]. Deterministic [312, 63, 155, 266]. deviation [106]. deviations [110]. diagnosability [203, 83]. diagnosis [211, 203]. diagnostic [256]. Diagonal [132]. diagonalization [198]. diagrams [214]. diameter [338]. Dichotomies [60]. dichotomy [17]. differential [350, 32, 331]. Difficulty [302]. diffusion [191]. digraphs [243]. dimensional [166]. dimensions [356]. Diophantine [150]. directed [11]. direction [56]. Discrete [156, 275]. disjoint [280, 134]. disjunctive [24]. disks [27]. disoriented [154]. Dispersion [245, 307]. distance [64, 273]. distances [289]. Distributed [46, 265, 62, 259, 72]. diversity [14, 61]. divide [146]. division [150, 164, 38]. divorcing [180]. domain [17, 309]. domination [69, 90]. Double [19, 29]. double-cut-and-joins [29]. DQN [78]. driven [239]. dual [63]. during [10]. Dynamic [232, 9, 333, 247, 222, 25, 228, 307]. Dynamical [218]. Dynamically [24]. dynamics [75, 86].

e-healthcare [7]. Eden [210]. Edge [337, 286, 3, 2, 151, 109]. edges [289]. Editorial [281, 282, 22, 44, 305, 314, 327, 343, 359, 57, 70, 80, 91, 101, 113, 124, 135, 144, 152, 158, 167, 176, 185, 195, 204, 215, 226, 235, 248, 261, 271, 283, 291]. Efficient [198, 288, 266, 10, 211, 330, 251]. EFX [329, 67, 224]. Egalitarian [66]. elections [163]. emptiness [229]. Enabling [36]. enclosure [105]. encoded [290]. encrypt [74]. encryption [123, 97]. end [318]. end-to-end [318]. energy [239, 103]. enhancement [74]. entangled [317]. entropy [221]. enumerating [288]. enumeration [349]. environments [233]. equal [73]. equal-letter [73]. equational [150]. equations [32]. equilibria [116]. equitable [175]. equivalence [311]. equivalent [24, 150]. erroneous [104]. error [47]. estimation [265, 132]. Eternal [225]. Euclidean [38]. EV [239]. evacuation [184, 156, 353]. Evaluating [109]. evolutionary [75, 246]. Exact [64, 65, 169, 132, 95, 37]. exchange [32]. execution [237]. Existence [202, 15, 224]. exits [156]. expansions [200]. expected [148]. explainable [279]. exploration [307, 287]. exponential [354]. expressibility [274]. expression [229]. Expressive [303, 332]. extendability [351]. extension [122]. Extensional [100]. extra [203, 109]. Extremal [5].

facilities [273]. facility [33, 331]. failure [50]. failures [323]. fair [269, 251, 164]. Fast [348, 151, 132, 95, 212, 77]. Faster [89]. Fault [50, 184, 192, 256, 109, 19]. Fault-tolerant [50, 184]. faults [189]. Feasibility [107, 219]. February [58, 71, 81, 92]. ferromagnetic [82]. few [333, 171]. Fibonacci [308]. fields [339]. Figure [187]. filling [296]. finder [77]. Finding [140, 251, 31, 9, 214, 12]. fine [59, 194]. fine-grained [59, 194]. Finer [59]. Finer-grained [59]. finite [341, 355, 225, 118]. Finiteness [342, 217]. first [267]. first-order [267]. fixed [325]. floorplan [202]. folding [296]. foldings [115]. forests [325]. formalization [32]. formation [118, 181]. Forming [270]. foundations [180]. FPT [79, 88]. fragments [267]. framework [304]. free [243, 87, 220, 120, 139, 133, 130, 123, 112, 88, 319, 121]. freshness [61]. frontier [329]. Full [309]. fully [117]. fully-decentralized [117]. functional [123, 309]. functions [277, 128, 194, 266]. funneled [15]. Further [295]. fusion [332]. fuzzy [241].

game [174, 241, 270, 223]. games

[240, 116, 149, 190]. Garden [210]. GaTEx [170]. Gathering [54, 316, 16]. gem [220]. Generalised [268]. generalized [90, 96]. generate [115]. generation [290, 222]. genomes [29]. geodesic [298]. geodetic [119]. Geography [21]. Geometric [339, 354, 137, 297]. given [219]. global [260]. good [83]. gradient [75]. grained [59, 194]. grammars [332]. graph [240, 121, 94, 232, 169, 346, 320, 335, 294, 36, 132, 147]. Graphs [319, 310, 82, 220, 120, 11, 156, 173, 139, 133, 130, 351, 225, 298, 188, 263, 119, 151, 157, 198, 218, 170, 245, 149, 338, 247, 175, 69, 6, 278, 307, 257, 90, 191, 98]. Gray [290, 115]. Greedy [88]. grid [225, 118, 312, 16, 181]. grids [20]. group [253, 164]. groups [342]. growth [354]. Grundy [295]. guarantees [269, 164, 234].

Hamiltonian [263]. handle [317]. handling [323]. hard [255, 121, 107, 170]. hardening [131]. Hardness [163, 59, 90]. Harmonious [352]. healthcare [7]. heavy [110]. heavy-tailed [110]. height [200]. helps [312]. hiding [7]. high [323, 36]. higher [166]. higher-dimensional [166]. Highly [94]. Holant [17]. homothetic [105]. horn [345]. hull [280]. hulls [280]. hybrid [238, 318]. hyperbolic [298]. hypercube [96, 111]. hypotheses [8].

Identity [97]. Identity-based [97]. Implementing [55]. Improved [193, 88, 78, 208, 234]. improves [183]. incentive [358]. inclusive [192]. increasing [242]. Incremental [229]. Independent [133, 118, 49, 170, 147]. independently [30]. index [5]. indivisible [67, 251]. induced [208]. inexact [65]. infinite [225, 20, 16]. influence [160]. information [42, 7]. information-based [42]. inhibition [222]. inhibitorless [260]. Injective [6]. inner [123]. Inoculation [157]. inputs [171]. insertions [26]. Instance [42]. Instance-optimal [42]. instances [121]. instantiation [97]. Integer [201]. Integer-valued [201]. integers [179]. integrity [43]. intensional [100]. interactions [86]. intersection [355, 229]. interval [300, 84]. intractability [329]. inversions [131]. IPC [142]. irregular [94]. isogeny [155]. isogeny-based [155]. isomorphism [294, 182]. isoperimetric [109]. Issue [197]. issues [69]. iterative [143, 340, 115].

January [23, 45]. joins [29]. Josephus [49, 109]. Journey [281]. July [216, 227, 236]. June [196, 205]. just [210].

kernel [208, 88]. Kernelization [330]. keys [155]. keyword [7]. Kirchhoff [5]. knapsack [336]. knowledge [13, 179, 182]. Kolmogorov [174]. Kotzig [21].

label [320]. ladder [349]. language [199]. languages [108]. Laplacian [132]. Large [106, 247, 31]. larger [254]. Largest [280]. lateral [222]. lattices [97]. layered [76]. LCAs [265]. learning [62, 143, 304, 39, 78, 160]. learning-based [304]. least [110]. left [276, 290]. left-child [290]. lenses [121]. letter [73]. level [258]. leveling [347]. Levenshtein [254]. Life [223]. limited [54]. Limiting [1, 41]. line [353, 287]. lineage [194]. Linear [129, 170, 233, 275, 172]. Linear-time [170]. link [109, 138]. linkable [141]. list [335]. load [9, 253]. local [211, 203, 161, 25, 46]. Locally [340, 278]. Locally-iterative [340]. Locating [273]. location [193, 33, 331]. locations [33]. logic [303, 299, 267, 228]. longest [242]. lose [174]. lotteries [349]. Loveland [174]. low [126]. Lower [207, 147, 72]. lower-bounded [72]. Lua [238].

machine [326, 39]. machines [230]. maintenance [247]. majority [86]. makespan [246]. Mapping [121]. March [102, 114, 125, 136]. martingales [201]. masking [131]. mass [55]. match [134]. Matching [67, 338, 63, 208]. Matching-based [67]. matchings [352, 162]. matchmaking [97]. Mathematical [238]. matrices [126, 198]. matrix [171]. matroid [324]. matroidal [36, 111]. max [75, 88, 99, 51]. Maximal [255, 288, 147]. maximin [164]. maximization [264, 160, 138]. maximizing [266]. Maximum [118, 247, 133, 230, 300, 12]. May [168, 177, 186]. meanders [115]. measures [231]. mechanism [264, 358]. median [331, 336]. medical [318]. meet [323, 312]. merge [276]. merged [242]. method [62, 160]. methods [32]. metric [72]. metrics [277]. min [154, 99]. min-max [99]. min-move [154]. minimal [146]. minimization [334, 61]. Minimizing [289, 326, 103]. Minimum [316, 337, 286, 302, 93, 3, 2, 27, 140, 273]. minor [60]. mixed [18]. mixed-charging [18]. mixing [116]. mobile [285, 316, 20, 245, 181, 358]. mobility [50]. modal [318]. Model [259, 358, 117, 68, 211, 46, 299, 127, 203, 141, 245, 19, 183, 83, 84]. model-based [211]. models [106, 238, 228]. Modified [1, 41, 83]. module [202]. modulo [34]. monadic [48]. Monitoring [119]. Monochromatic [308]. Monodic [267]. monographs [40]. monoid [217]. monotone [171]. morphisms [231]. Morse [308]. most [278]. Move [181, 154]. Move-optimal [181]. multi [348, 75, 320, 9, 253, 108, 300, 318, 7]. multi-agent [75]. multi-album [348]. multi-cover [300]. multi-keyword [7]. multi-label [320]. multi-load [9, 253]. multi-modal [318]. multi-pattern [108]. MultiCover [37]. Multidepot [234]. Multiple [222, 156, 256]. multiplication [34]. multiprocessor [256, 83]. multivalued [189]. mutual [154, 50]. myopic [118, 16].

Nash [116]. Near [324, 181]. Near-optimal [324, 181]. negative [277]. neighbor [83].

Network [149, 192, 96]. networks

[24, 18, 193, 346, 15, 218, 203, 178, 28, 36, 19, 25, 212, 180, 134, 10, 138]. Neumann [221]. neural [222]. Newton [62]. Nim [21]. node [19]. nodes [325]. noisy [86]. non [229, 192, 140, 112, 275]. non-adaptive [112]. non-dense [275]. non-emptiness [229]. non-inclusive [192]. non-terminal [140]. nondeterministic [199]. note [146]. Notes [161]. notice [337, 286, 41]. novel [256]. NP [255, 170]. NP-hard [255, 170]. number [73, 173, 325]. numbers [244, 150, 225, 294]. numerical [32]. Nyldon [268].

objects [194]. oblivious [20, 54]. obstacle [287]. occurrences [95]. October [328, 344, 360]. On-line [287]. one [150, 334, 287]. one-one [150]. one-sided [334]. online [239, 63, 326]. Ono [150]. opaqueness [165]. open [174]. Opinion [191, 138, 86]. optimal [193, 42, 116, 16, 109, 349, 181, 214, 224, 324]. optimization [297, 324]. Optimizing [143, 1, 41]. Oracle [112]. order [275, 267]. ordering [302, 214]. orderings [170]. Ordinal [164]. organic [127]. orientations [15]. oriented [119, 358]. orienteering [11]. outerplanar [294]. outliers [79]. output [265]. overlaps [47]. overlays [323].

P [222]. Packing [30, 335]. pairing [123]. pairing-free [123]. Papers [197]. parallel [212, 32, 345]. parallelogram [169]. parameterizations [43]. Parameterized [93, 320, 322, 336, 295, 208, 89]. Parameterizing [85]. parameters [60, 175]. parent [180]. Pareto [224]. Pareto-optimal [224]. parity [212, 321]. parity-constrained [321]. Parking [20]. Partial [258, 4, 32]. participatory [301]. partition [330]. partitioning [103]. partitions [85, 28, 98]. partizan [240]. Path [28, 307, 193, 263, 85, 52, 9, 134, 98]. paths [244, 188]. pattern [192, 106, 181, 108]. Payment [84]. pbn [304]. pbn-STAC [304]. perfect [120, 170]. performance [246, 183]. perimeter [233]. period [10]. periodic [320]. permutation [263]. Personal [281]. Phase [86]. phylogenetic [15, 28]. pickup [9, 253]. pictures [311]. placements [53]. planar [173, 53]. planarity [258]. plane [296]. plane-filling [296]. planning [356]. played [149]. ply [27]. PMC [211, 203]. Point [105, 53]. points [27]. policy [75]. polygons [105]. polylogarithmic [172]. polynomial [99]. polynomials [77]. polyphenyl [5]. Popularity [14]. population [35]. Post [155]. Post-quantum [155]. power [165, 332]. Powers [126]. precision [172]. prefix [29]. preimages [223]. presence [189, 50]. primal [63]. primal-dual [63]. prime [309]. primes [357]. primitive [106]. priority [8]. Privacy [35, 331]. Private [110]. probabilistic [35, 267, 199]. probe [356]. Problem [337, 286, 105, 350, 87, 37, 355, 154, 11, 20, 206, 3, 2, 64, 116, 14, 239, 254, 52, 242, 9, 253, 109, 137, 182, 69, 99, 95, 342, 246, 264, 321, 336, 10]. problems [316, 347, 133, 170, 26, 297, 326, 90]. product [123]. products [171]. programming [333, 129]. programs [207, 259]. progressions [308]. projection [299]. proofs [250]. properties [166, 106, 40]. proportional [76]. propositional [299]. protocols [35, 184]. pruning [62]. pruning-based [62]. pseudoforest [89]. pseudoinverse [132]. PSPACE [107]. PSPACE-hard [107]. punctual [275].

quadruple [69]. quality [228]. Quantum [334, 214, 333, 355, 221, 112, 155, 146]. qubits [333]. queries [147].

R [353]. radius [338]. Ramsey [244]. Random [200, 325, 339]. randomly [163]. randomness [221]. rank [126, 190]. rank-based [190]. ranking [294]. Rate [1, 41, 219]. rates [324]. ratio [169]. rational [150, 106]. reactantless [260]. reaction [260]. realization [320]. recommendation [138]. Reconfigurable [178]. Reconfiguration [349]. reconstruction [254, 147]. reconstructional [274]. rectangle [76]. rectangular [181]. recurrent [107]. Recursive [115, 134]. red [18, 137]. red-black [18]. reductions [59, 201]. refinement [274, 213, 241]. reflection [122]. refutability [345]. region [287]. regret [61]. regular [229, 128]. reinforcement [304, 78]. related [316, 133]. relation [341]. relations [241]. reliability [192, 346, 36, 109, 19]. reliable [278]. removal [222]. rendezvous [312]. Repetitiveness [231]. representations [76]. reprogramming [304]. requirement [273]. rerandomized [155]. resource [347, 143]. respect [5]. responses [25]. Restricted [17, 232, 252]. Results [319, 347, 295, 254, 132, 175]. Retention [206]. Retraction [337, 286, 41]. reversible [87]. revisited [351, 100]. rewriting [87]. ring [141, 309]. robotic [356, 219]. robots [285, 316, 154, 20, 118, 54, 165, 16, 353, 245, 181]. Robust [323, 324]. robustness [213]. Roman [69]. roommate [66, 14]. root [77]. root-finder [77]. rooted [15]. rotation [115, 146]. rounds [340]. routing [178, 234]. routings [134]. Rudin [308]. rules [222, 279]. rumor [10]. runs [73]. Runtime [246].

S [353]. S/R [353]. safety [10]. salesman [350]. Salomaa [187]. Sampling [265, 325]. SARSA [78]. SAT [299, 88, 51, 95]. SAT-based [299]. Scalability [1, 41]. Scale [1, 41]. scenario [4]. scenarios [52]. scheduling [35, 84, 239, 326, 219, 246]. scheme [139, 141, 88, 182, 7]. Sci [3]. Science [187, 286, 337, 41]. search [65, 7]. secure [155]. Securing [123]. security [74, 97]. segments [297]. Selected [197, 163]. selection [61, 322, 358]. Self [68, 131, 189, 294]. self-delimiting [294]. Self-masking [131]. Self-similarity [68]. Self-stabilizing [189]. semantics [267, 237, 199]. semi [115, 326]. semi-meanders [115]. semi-online [326]. sensing [318]. sensitive [252, 7]. separation [303]. separations [112]. September [284, 292, 306, 315]. sequence [254]. sequences [290]. Sequential [160]. service [228]. Services [1, 41]. Set [270, 53, 118, 4, 357, 140, 137, 147, 322, 134]. set-to-set [134]. sets [133, 119, 61, 170, 174, 39]. settings [75]. SGD [117]. shaped [202]. shapes [354]. Shapiro [308]. sharing [65, 190, 7]. Shell [148]. shift [311]. shift-equivalence [311]. short [279]. shortest [188, 52]. should [213]. side [65]. side-sharing [65]. sided [334]. Sign [74]. Sign-then-encrypt [74]. signals [318]. signature [141, 182]. signatures [155]. signed [218]. Similarity [277, 68]. Simple [325, 188, 151]. simulations [348]. single [326]. sink [193]. size [207, 123, 250, 280]. size-change [250]. sizes [316, 254, 17]. small [207, 198, 140, 31]. small-size [207]. Smyth [161]. Smyth-completes [161]. Sniffing [312]. social [10, 138]. soft [75]. soft-max [75]. Software [1, 41]. solution [109]. solutions [212]. Solving [170, 229, 32]. Some [175, 319, 297]. sort [148, 19]. Sorting [29, 294]. space [72]. spanners [18, 289]. Spanning [121, 337, 286, 169, 3, 2, 140, 49, 247, 96]. sparse [126, 335]. spatial [181]. Special [197]. specified [55]. speed [55]. speedup [333]. spending [269]. sphere [12]. spiking [222]. spin [82]. split [278]. square [273]. squares [27, 30]. stability [66, 218]. stabilizing [189]. Stable [217]. STAC [304]. stamp [115]. standard [141]. star [192]. state [183]. statistics [106]. stimuli [25]. stochastic [10]. storage [303]. strategic [313]. strategies [157, 174]. strikes [88]. string [231, 146]. strings [288]. strong [61]. stronger [97]. strongly [99]. Structural [43, 60]. Structure [223, 211]. structured [194]. structures [241]. study [307]. subclasses [120, 6]. Subgraph [337, 286, 3, 2, 182]. sublinear [340]. Submodular [301, 266]. Subnetwork [346, 19]. subquadratic [139]. Subsequence [209, 242]. subsequences [288, 171]. Subset [37, 130]. sum [93, 207, 302]. Super [96]. Suppes [150]. supplier [321]. surjectivity [210]. surjunctivity [217]. sweep [252]. symbolic [237]. symmetric [198, 172, 266]. synchronisation [212]. synchronous [259]. Synergistic [13]. systems [341, 55, 260, 82, 303, 256, 222, 104, 296, 83, 345, 7].

tables [238]. tailed [110]. takes [210]. tandems [65]. target [322]. targets [252]. Task [253]. tasks [107, 103, 358]. TCS [281, 282, 197]. Telephone [257]. temporal [352, 299, 320, 267, 228]. tensors [172]. terminal [140]. Termination [87, 250]. terms [192]. tests [317]. Theor [3]. Theoretical [187, 337, 286, 41, 180]. theory [48, 194, 100]. thief [11]. three [356]. threshold [218]. throughput [230, 219]. Thue [308]. Tiling [311]. Time [16, 239, 252, 170, 172, 326, 99]. time-sensitive [252]. Tolerance [34]. tolerant [184, 50]. toric [48]. toroidal [311]. total [326]. tournaments [31]. Towering [187]. tractable [121]. training [358]. trajectory [356]. transducer [128]. transform [73]. transformations [180, 40]. transformer [318]. transition [86]. transitive [243]. transitive-free [243]. translocation [64]. TransMCS [318]. Transparency [165]. transposition [346]. traveling [350]. tree [289, 60, 140, 247, 175, 121]. tree-coloring [175]. trees [325, 49, 290, 200, 279, 8]. treewidth [188, 198, 257]. triangular [16]. Tribute [187]. Truthful [33]. TSP [333]. Turing [317, 201]. two [329, 288, 33, 273, 257]. two-facility [33]. type [100].

UAV [252]. UAV-based [252]. unbounded [123, 287]. unconditional [141]. undecidability [311]. undirected [218]. Unfolding [183]. Uniqueness [311]. Unit [345, 27]. universal [255]. Unpaired [134]. unrelated [230]. unsolvability [150]. updating [213]. used [32]. Using [1, 41, 123, 259, 212, 339, 181, 309]. utilities [190].

validity [150]. valued [201]. variable [214]. variables [95, 183]. variants [21, 338]. varietal [111]. vector [123, 171]. vehicle [234]. vehicular [18]. Verifiable [7]. Vertex [220, 49, 93, 130, 225, 43, 302]. Vertex-critical [220]. Vertex-independent [49]. vertices [289, 53]. very [333]. via [36, 122, 138]. video [317]. video-based [317]. Visibility [122, 154, 54, 50]. visible [252]. voters [163]. voting [42, 313]. vs [165].

wait [213]. wallets [155]. way [63]. weak [276, 295]. weight [133, 88]. Weighted [162, 230, 322]. welfare [264]. Wheeler [73]. width [255]. win [293]. winner [313]. within [202]. without [50]. words [47, 48, 209, 268, 308]. Worker [358]. Worst [269]. Worst-case [269].

Yoneda [161]. Yoneda-completes [161].

zero [179, 182]. zero-knowledge [179].

References

Latha:2023:OSA

[1] V. L. Padma Latha, N. Sudhakar Reddy, and A. Suresh Babu. Optimizing scalability and availability of cloud based software services using modified scale rate limiting algorithm. Theoretical Computer Science, 943(??): 230–240, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                             //www.sciencedirect.com/science/
                                             article/pii/S0304397522004418. See

retraction notice [41].

Civril:2023:NAA

[2] A. Çivril. A new approximation algorithm for the minimum 2-edge-connected spanning subgraph problem. Theoretical Computer Science, 943(??): 121–130, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                             //www.sciencedirect.com/science/
                                             article/pii/S0304397522007368. See

corrigendum [3] and retraction notice [286].

Civril:2023:CNA

[3] A. Çivril. Corrigendum to “A new approximation algorithm for the minimum 2-edge-connected spanning subgraph problem” [Theor. Comput. Sci. 943 (2023) 121–130]. Theoretical Computer Science, 963(??):??, June 26, 2023. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                             //www.sciencedirect.com/science/
                                             article/pii/S0304397523002396. See

Dimant:2025:APS

[4] Shai Michael Dimant and Sven O. Krumke. On approximating partial scenario set cover. Theoretical Computer Science, 1023(??):??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005085.

Li:2025:EPC

[5] Chengmin Li, Hong Bian, and Haizheng Yu. Extremal polyphenyl chains with respect to the Kirchhoff index. Theoretical Computer Science, 1023 (??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005103.

Panda:2025:ICS

[6] B. S. Panda and Rumki Ghosh. Injective coloring of subclasses of chordal graphs. Theoretical Computer Science, 1023(??):??, January 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005115.

Zhao:2025:VAB

[7] Jie Zhao, Hejiao Huang, Yongliang Xu, Xiaojun Zhang, Hongwei Du, and Chao Huang. Verifiable attribute-based multi-keyword search scheme with sensitive information hiding for cloudassisted e-healthcare sharing systems. Theoretical Computer Science, 1023 (??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005127.

Yuan:2025:ADT

[8] Jing Yuan and Shaojie Tang. Approximating decision trees with priority hypotheses. Theoretical Computer Science, 1023(??):??, January 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005139.

Li:2025:DPF

[9] Yifei Li, Ruixi Huang, Hao Ye, Hejiao Huang, and Hongwei Du. Dynamic path finding for multi-load agent pickup and delivery problem. Theoretical Computer Science, 1023 (??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005140.

Zhu:2025:EAS

[10] Jianming Zhu, Ye Xing, Runzhi Li, Smita Ghosh, Priyanshi Garg, and Weili Wu. Efficient algorithm for stochastic rumor blocking problem in social networks during safety accident period. Theoretical Computer Science, 1023(??):??, January 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005152.

Bloch-Hansen:2025:ATO

[11] Andrew Bloch-Hansen, Roberto Solis-Oba, and Daniel R. Page. Algorithms for the thief orienteering problem on directed acyclic graphs. Theoretical Computer Science, 1023 (??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005176.

Zhang:2025:AAF

[12] Kaiqi Zhang, Siyuan Zhang, Jirun Gao, Hongzhi Wang, Hong Gao, and Jianzhong Li. Approximation algorithms for finding maximum containing circle and sphere. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005188.

Cachin:2025:SK

[13] Christian Cachin, David Lehnherr, and Thomas Studer. Synergistic knowledge. Theoretical Computer Science, 1023(??):??, January 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752400519X.

Ge:2025:PRD

[14] Steven Ge and Toshiya Itoh. Popularity on the roommate diversity problem. Theoretical Computer Science, 1023(??):??, January 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005206.

Docker:2025:EFO

[15] Janosch Döcker and Simone Linz. On the existence of funneled orientations for classes of rooted phylogenetic networks. Theoretical Computer Science, 1023(??):??, January 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005255.

Goswami:2025:TOG

[16] Pritam Goswami, Avisek Sharma, Satakshi Ghosh, and Buddhadeb Sau. Time optimal gathering of myopic robots on an infinite triangular grid. Theoretical Computer Science, 1023 (??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005474.

Liu:2025:RHD

[17] Yin Liu, Austen Z. Fan, and Jin-Yi Cai. Restricted Holant dichotomy on domain sizes 3 and 4. Theoretical Computer Science, 1023(??):??, January 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005486.

Bereg:2025:CRB

[18] Sergey Bereg, Yuya Higashikawa, Naoki Katoh, Junichi Teruyama, Yuki Tokuni, and Binhai Zhu. Constructing red-black spanners for mixed-charging vehicular networks. Theoretical Computer Science, 1023(??):??, January 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005498.

Meng:2025:DDS

[19] Kaiyue Meng and Yuxing Yang. Double declined subnetwork reliability analysis in bubble-sort networks under node fault model. Theoretical Computer Science, 1023(??):??, January 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005589.

Chakraborty:2025:PPO

[20] Abhinav Chakraborty and Krishnendu Mukhopadhyaya. Parking problem by oblivious mobile robots in infinite grids. Theoretical Computer Science, 1023(??):??, January 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752400570X.

Arun:2025:GKN

[21] Srinivas Arun. Geography, Kotzig’s Nim, and variants. Theoretical Computer Science, 1023(??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005747.

Anonymous:2025:EBa

[22] Anonymous. Editorial Board. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524006091.

Anonymous:2025:Ja

[23] Anonymous. 1 January 2025. Theoretical Computer Science, 1023(??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aracena:2025:DED

[24] Julio Aracena, Luis Cabrera-Crot, Adrien Richard, and Lilian Salinas. Dynamically equivalent disjunctive networks. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005164.

Oh:2025:ACR

[25] Shunhao Oh, Dana Randall, and Andréa W. Richa. Adaptive collective responses to local stimuli in anonymous dynamic networks. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005218.

Ibarra:2025:DPC

[26] Oscar H. Ibarra and Ian McQuillan. On decision problems concerning contextual insertions and deletions. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752400522X.

Durocher:2025:AAM

[27] Stephane Durocher, J. Mark Keil, and Debajyoti Mondal. Approximation algorithms for minimum ply covering of points with unit squares and unit disks. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005231.

Lafond:2025:PPP

[28] Manuel Lafond and Vincent Moulton. Path partitions of phylogenetic networks. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005243.

Fertin:2025:SGP

[29] Guillaume Fertin, Géraldine Jean, and Anthony Labarre. Sorting genomes by prefix double-cut-and-joins. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005267.

Wu:2025:PSI

[30] Wei Wu, Hiroki Numaguchi, Nir Halman, Yannan Hu, and Mutsunori Yagiura. Packing squares independently. Theoretical Computer Science, 1024 (??):??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005279.

Yuster:2025:FCS

[31] Raphael Yuster. Finding and counting small tournaments in large tournaments. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005280.

Samaniego:2025:FPD

[32] Cristóbal Samaniego and Guillaume Houzeaux. A formalization of parallel data exchange algorithms used by numerical methods for solving partial differential equations. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005292.

Kanellopoulos:2025:TTF

[33] Panagiotis Kanellopoulos, Alexandros A. Voudouris, and Rongsen Zhang. Truthful two-facility location with candidate locations. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005309.

Gupta:2025:TAA

[34] Arya Tanmay Gupta and Sandeep S. Kulkarni. Tolerance to asynchrony in algorithms for multiplication and modulo. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005310.

Amir:2025:PPP

[35] Talley Amir and James Aspnes. Privacy in population protocols with probabilistic scheduling. Theoretical Computer Science, 1024(??):??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005437.

Li:2025:EHR

[36] Xiao-Yan Li, Zhaoding Lin, Hongbin Zhuang, and Jou-Ming Chang. Enabling high reliability via matroidal connectivity and conditional matroidal connectivity on arrangement graph networks. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005449.

Benoist:2025:ESM

[37] Emile Benoist, Guillaume Fertin, and Géraldine Jean. The Exact Subset MultiCover problem. Theoretical Computer Science, 1024(??):??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752400553X.

Pin:2025:EDB

[38] Jean-Éric Pin. Euclidean division by d in base b. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005541.

Singh:2025:CML

[39] Barkha Singh, Sreedevi Indu, and Sudipta Majumdar. Comparison of machine learning algorithms for classification of Big Data sets. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005553.

delaTour:2025:APT

[40] Thierry Boy de la Tour. Algebraic properties and transformations of monographs. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005565.

Latha:2025:RNO

[41] V. L. Padma Latha, N. Sudhakar Reddy, and A. Suresh Babu. Retraction notice to “Optimizing Scalability and Availability of Cloud Based Software Services Using Modified Scale Rate Limiting Algorithm” [Theoretical computer science 943 (2023) 230– 240]. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005619. See

[1].

Chierichetti:2025:IOI

[42] Flavio Chierichetti. Instance-optimal information-based voting. Theoretical Computer Science, 1024(??):??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005620.

Gima:2025:SPV

[43] Tatsuya Gima, Tesshu Hanaka, Yasuaki Kobayashi, Ryota Murai, Hirotaka Ono, and Yota Otachi. Structural parameterizations of vertex integrity. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005711.

Anonymous:2025:EBb

[44] Anonymous. Editorial Board. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524006170.

Anonymous:2025:Jb

[45] Anonymous. 12 January 2025. Theoretical Computer Science, 1024(??):??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delporte-Gallet:2025:DCA

[46] Carole Delporte-Gallet, Hugues Fauconnier, Pierre Fraigniaud, and Mikaël Rabie. Distributed computing in the asynchronous LOCAL model. Theoretical Computer Science, 1025(??): ??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005693.

Anselmo:2025:DAW

[47] Marcella Anselmo, Manuela Flores, and Maria Madonia. Density of kary words with 0, 1, 2-error overlaps. Theoretical Computer Science, 1025 (??):??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005759.

Berthe:2025:MTT

[48] Valérie Berthé, Toghrul Karimov, Joris Nieuwveld, Joël Ouaknine, Mihir Vahanwala, and James Worrell. The monadic theory of toric words. Theoretical Computer Science, 1025(??): ??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005760.

He:2025:VIS

[49] Qi He, Yan Wang, Jianxi Fan, and Baolei Cheng. Vertex-independent spanning trees in complete Josephus cubes. Theoretical Computer Science, 1025 (??):??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005863.

Pramanick:2025:FTM

[50] Subhajit Pramanick, Saswata Jana, and Partha Sarathi Mandal. Faulttolerant mutual visibility without any axis agreement in presence of mobility failure. Theoretical Computer Science, 1025(??):??, February 2, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005875.

Panolan:2025:MSC

[51] Fahad Panolan and Hannane Yaghoubizade. On MAX–SAT with cardinality constraint. Theoretical Computer Science, 1025(??):??, February 2, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005887.

Kasperski:2025:ASP

[52] Adam Kasperski and Paweł Zieliński. Approximating the shortest path problem with scenarios. Theoretical Computer Science, 1025(??):??, February 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005899.

Damaschke:2025:CPN

[53] Peter Damaschke, Fredrik Ekstedt, and Raad Salman. On central placements of new vertices in a planar point set. Theoretical Computer Science, 1025 (??):??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005905.

DiLuna:2025:GCL

[54] Giuseppe Antonio Di Luna, Ryuhei Uehara, Giovanni Viglietta, and Yukiko Yamauchi. Gathering on a circle with limited visibility by anonymous oblivious robots. Theoretical Computer Science, 1025(??):??, February 2, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005917.

Anderson:2025:CMA

[55] David F. Anderson and Badal Joshi. Chemical mass–action systems as analog computers: Implementing arithmetic computations at specified speed. Theoretical Computer Science, 1025 (??):??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006005.

Bensmail:2025:ADC

[56] Julien Bensmail, Hervé Hocquard, and Clara Marcille. Adding direction constraints to the 1–2–3 Conjecture. Theoretical Computer Science, 1025 (??):??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006029.

Anonymous:2025:EBc

[57] Anonymous. Editorial Board. Theoretical Computer Science, 1025(??): ??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006340.

Anonymous:2025:Fa

[58] Anonymous. 2 February 2025. Theoretical Computer Science, 1025(??):??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abboud:2025:FGR

[59] Elie Abboud and Noga Ron-Zewi. Finer-grained reductions in fine-grained hardness of approximation. Theoretical Computer Science, 1026(??): ??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524005930.

Gima:2025:DTM

[60] Tatsuya Gima, Soh Kumabe, Kazuhiro Kurita, Yuto Okada, and Yota Otachi. Dichotomies for tree minor containment with structural parameters. Theoretical Computer Science, 1026(??): ??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006017.

Guo:2025:TSR

[61] Hongjie Guo, Jianzhong Li, and Hong Gao. Towards strong regret minimization sets: Balancing freshness and diversity in data selection. Theoretical Computer Science, 1026(??): ??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006030.

Chen:2025:APB

[62] Shuzhen Chen, Yuan Yuan, Youming Tao, Tianzhu Wang, Zhipeng Cai, and Dongxiao Yu. Adaptive pruning-based Newton’s method for distributed learning. Theoretical Computer Science, 1026(??):??, February 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006042.

Kakimura:2025:DPD

[63] Naonori Kakimura and Tomohiro Nakayoshi. Deterministic primaldual algorithms for online k-way matching with delays. Theoretical Computer Science, 1026(??):??, February 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006054.

Constantin:2025:EAA

[64] Maria Constantin and Alexandru Popa. Exact and approximation algorithms for the contiguous translocation distance problem. Theoretical Computer Science, 1026(??):??, February 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006200.

Marcus:2025:EIS

[65] Shoshana Marcus, Dina Sokol, and Sarah Zelikovitz. Exact and inexact search for 2d side-sharing tandems. Theoretical Computer Science, 1026 (??):??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006224.

Bonifaci:2025:ERA

[66] Vincenzo Bonifaci and Helena Rivera Dallorto. Egalitarian roommate allocations: Complexity and stability. Theoretical Computer Science, 1026 (??):??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006261.

Kobayashi:2025:EAI

[67] Yusuke Kobayashi, Ryoga Mahara, and Souta Sakamoto. EFX allocations for indivisible chores: Matching-based approach. Theoretical Computer Science, 1026(??):??, February 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006273.

Barrett:2025:SSC

[68] Jordan Barrett, Bogumił Kamiński, Paweł Prałat, and François Théberge. Self-similarity of communities of the ABCD model. Theoretical Computer Science, 1026(??):??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0304397524006297.

Palagiri:2025:CIC

[69] Venkata Subba Reddy Palagiri, Guru Pratap Sharma, and Ismael G. Yero. Complexity issues concerning the quadruple Roman domination problem in graphs. Theoretical Computer Science, 1026 (??):??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006303.

Anonymous:2025:EBd

[70] Anonymous. Editorial Board. Theoretical Computer Science, 1026(??): ??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006595.

Anonymous:2025:Fb

[71] Anonymous. 12 February 2025. Theoretical Computer Science, 1026(??):??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liang:2025:DAL

[72] Ting Liang, Xiaoliang Wu, Jinhui Xu, and Qilong Feng. The distributed algorithms for the lower-bounded kcenter clustering in metric space. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524005929.

Biagi:2025:NEL

[73] Elena Biagi, Davide Cenzato, Zsuzsanna Lipták, and Giuseppe Romana. On the number of equal-letter runs of the bijective Burrows–Wheeler transform. Theoretical Computer Science, 1027(??):??, February 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006212.

Chen:2025:STE

[74] Yanbo Chen and Yunlei Zhao. Signthen-encrypt with security enhancement and compressed ciphertext. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006236.

Bernasconi:2025:EDS

[75] Martino Bernasconi, Federico Cacciamani, Simone Fioravanti, Nicola Gatti, and Francesco Trovò. The evolutionary dynamics of soft-max policy gradient in multi-agent settings. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006285.

Haase:2025:LAP

[76] Carolina Haase and Philipp Kindermann. On layered area-proportional rectangle contact representations. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006388.

Pan:2025:NFR

[77] Victor Y. Pan, Soo Go, Qi Luan, and Liang Zhao. A new fast rootfinder for black box polynomials. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752400639X.

Yao:2025:ISD

[78] Guangyu Yao, Nan Zhang, Zhenhua Duan, and Cong Tian. Improved SARSA and DQN algorithms for reinforcement learning. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752400642X.

Dabas:2025:FAC

[79] Rajni Dabas, Neelima Gupta, and Tanmay Inamdar. FPT approximation for capacitated clustering with outliers. Theoretical Computer Science, 1027(??):??, February 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006431.

Anonymous:2025:EBe

[80] Anonymous. Editorial Board. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006765.

Anonymous:2025:Fc

[81] Anonymous. 19 February 2025. Theoretical Computer Science, 1027(??):??, February 19, 2025. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Bai:2025:CFS

[82] Zonglei Bai, Yongzhi Cao, and Hanpin Wang. The complexity of ferromagnetic 2-spin systems on bounded degree graphs. Theoretical Computer Science, 1028(??):??, February 28, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524005577.

Wang:2025:GND

[83] Mu-Jiang-Shan Wang, Dong Xiang, and Sun-Yuan Hsieh. G-good-neighbor diagnosability under the modified comparison model for multiprocessor systems. Theoretical Computer Science, 1028(??):??, February 28, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006443.

Friedetzky:2025:PSI

[84] Tom Friedetzky, David C. Kutner, George B. Mertzios, Iain A. Stewart, and Amitabh Trehan. Payment scheduling in the Interval Debt Model. Theoretical Computer Science, 1028 (??):??, February 28, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006455.

Fernau:2025:PPP

[85] Henning Fernau, Florent Foucaud, Kevin Mann, Utkarsh Padariya, and Rajath Rao K. N. Parameterizing path partitions. Theoretical Computer Science, 1028(??):??, February 28, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006467.

dAmore:2025:PTM

[86] Francesco d’Amore and Isabella Ziccardi. Phase transition of the 3-majority opinion dynamics with noisy interactions. Theoretical Computer Science, 1028(??):??, February 28, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006479.

Barile:2025:TRR

[87] Adriano Barile, Stefano Berardi, and Luca Roversi. Termination of rewriting on reversible Boolean circuits as a free 3-category problem. Theoretical Computer Science, 1028(??): ??, February 28, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006480.

Manurangsi:2025:IFA

[88] Pasin Manurangsi. Improved FPT approximation scheme and approximate kernel for biclique-free max kweight SAT: Greedy strikes back. Theoretical Computer Science, 1028(??): ??, February 28, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006509.

Tsur:2025:FPA

[89] Dekel Tsur. Faster parameterized algorithm for r-pseudoforest deletion. Theoretical Computer Science, 1028 (??):??, February 28, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006510.

Wang:2025:AHD

[90] Po Yuan Wang, Naoki Kitamura, Taisuke Izumi, and Toshimitsu Masuzawa. Approximation hardness of domination problems on generalized convex graphs. Theoretical Computer Science, 1028(??):??, February 28, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006522.

Anonymous:2025:EBf

[91] Anonymous. Editorial Board. Theoretical Computer Science, 1028(??): ??, February 28, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000064.

Anonymous:2025:Fd

[92] Anonymous. 28 February 2025. Theoretical Computer Science, 1028(??):??, February 28, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aute:2025:PAM

[93] Shubhada Aute and Fahad Panolan. Parameterized algorithms for minimum sum vertex cover. Theoretical Computer Science, 1029(??): ??, March 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006492.

Bensmail:2025:HIG

[94] Julien Bensmail, Malory Marin, Leandro Montero, and Alexandre Talon. Highly irregular graph decompositions. Theoretical Computer Science, 1029 (??):??, March 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006534.

Peng:2025:FEA

[95] Junqiang Peng and Mingyu Xiao. Fast exact algorithms for the SAT problem with bounded occurrences of variables. Theoretical Computer Science, 1029(??):??, March 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006546.

Wang:2025:SSC

[96] Xiaoqian Wang and Eminjan Sabir. Super spanning connectivity of the generalized hypercube network. Theoretical Computer Science, 1029(??): ??, March 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006558.

Wang:2025:IBM

[97] Yuejun Wang, Baocang Wang, Qiqi Lai, and Yu Zhan. Identity-based matchmaking encryption with stronger security and instantiation on lattices. Theoretical Computer Science, 1029 (??):??, March 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006650.

Zhao:2025:AAC

[98] Jingyang Zhao and Mingyu Xiao. Approximation algorithms for cycle and path partitions in complete graphs. Theoretical Computer Science, 1029 (??):??, March 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006662.

Pan:2025:SPT

[99] Pengxiang Pan and Hongtao Zhu. A strongly polynomial time approximation algorithm for the min-max clustered cycle cover problem. Theoretical Computer Science, 1029(??): ??, March 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006674.

Kapulkin:2025:ECI

[100] Krzysztof Kapulkin and Yufeng Li.

Extensional concepts in intensional type theory, revisited. Theoretical Computer Science, 1029(??):??, March 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006686.

Anonymous:2025:EBg

[101] Anonymous. Editorial Board. Theoretical Computer Science, 1029(??): ??, March 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000362.

Anonymous:2025:Ma

[102] Anonymous. 2 March 2025. Theoretical Computer Science, 1029(??):??, March 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2025:CME

[103] Wei Liu, Jian-Jia Chen, and Yongjie Yang. On the complexity of minimizing energy consumption of partitioning DAG tasks. Theoretical Computer Science, 1030(??):??, March 13, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006649.

Pittou:2025:ABA

[104] Maria Pittou and George Rahonis. An automata-based approach for addressing erroneous behaviors and deadlocks in component-based systems. Theoretical Computer Science, 1030 (??):??, March 13, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006698.

Akram:2025:PEP

[105] Waseem Akram and Sanjeev Saxena.

Point enclosure problem for homothetic polygons. Theoretical Computer Science, 1030(??):??, March 13, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006716.

Goldwurm:2025:LDP

[106] Massimiliano Goldwurm and Marco Vignati. Large deviation properties for pattern statistics in primitive rational models. Theoretical Computer Science, 1030(??):??, March 13, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006728.

Bonifaci:2025:FAR

[107] Vincenzo Bonifaci and Alberto Marchetti-Spaccamela. Feasibility analysis of recurrent DAG tasks is PSPACE-hard. Theoretical Computer Science, 1030 (??):??, March 13, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006790.

Xie:2025:CDC

[108] Jingnan Xie, Harry B. Hunt, and Richard E. Stearns. On the computational and descriptional complexity of multi-pattern languages. Theoretical Computer Science, 1030(??): ??, March 13, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000015.

Liu:2025:ERC

[109] Sufang Liu, Zhaoman Huang, Yueke Lv, and Chia-Wei Lee. Evaluating the reliability of complete Josephus cubes under extra link fault with the optimal solution of the edge isoperimetric problem. Theoretical Computer Science, 1030(??):??, March 13, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000027.

Wang:2025:PLA

[110] Di Wang and Jinhui Xu. Private least absolute deviations with heavy-tailed data. Theoretical Computer Science, 1030(??):??, March 13, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752500009X.

Yang:2025:CMC

[111] Lulu Yang, Shuming Zhou, Qifan Zhang, and Guanqin Lian. The (conditional) matroidal connectivity of varietal hypercube. Theoretical Computer Science, 1030(??):??, March 13, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000143.

Hepp:2025:OSN

[112] Henrique Hepp, Murilo V. G. da Silva, and Leandro M. Zatesko. Oracle separations for non-adaptive collapsefree quantum computing. Theoretical Computer Science, 1030(??): ??, March 13, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000167.

Anonymous:2025:EBh

[113] Anonymous. Editorial Board. Theoretical Computer Science, 1030(??): ??, March 13, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000465.

Anonymous:2025:Mb

[114] Anonymous. 13 March 2025. Theoretical Computer Science, 1030(??):??, March 13, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2025:RIA

[115] Bowie Liu, Dennis Wong, Chan-Tong Lam, and Marcus Im. Recursive and iterative approaches to generate rotation Gray codes for stamp foldings and semimeanders. Theoretical Computer Science, 1031(??):??, March 21, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397524006704.

Deng:2025:OMP

[116] Xiaotie Deng, Dongchen Li, and Hanyu Li. On the optimal mixing problem of approximate Nash equilibria in bimatrix games. Theoretical Computer Science, 1031(??):??, March 21, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000106.

Attiya:2025:AFD

[117] Hagit Attiya and Noa Schiller. Asynchronous fully-decentralized SGD in the cluster-based model. Theoretical Computer Science, 1031(??):??, March 21, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000118.

Das:2025:MIS

[118] Raja Das, Avisek Sharma, and Buddhadeb Sau. Maximum independent set formation on a finite grid by myopic robots. Theoretical Computer Science, 1031(??):??, March 21, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000155.

Das:2025:MAG

[119] Tapas Das, Florent Foucaud, Clara Marcille, P. D. Pavan, and Sagnik Sen. Monitoring arc-geodetic sets of oriented graphs. Theoretical Computer Science, 1031(??):??, March 21, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000179.

Bhyravarapu:2025:CFC

[120] Sriram Bhyravarapu, Subrahmanyam Kalyanasundaram, and Rogers Mathew. Conflict-free coloring on subclasses of perfect graphs and bipartite graphs. Theoretical Computer Science, 1031 (??):??, March 21, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000180.

Barros:2025:CFS

[121] Bruno José S. Barros, Luiz Satoru Ochi, Rian Gabriel S. Pinheiro, and Uéverton S. Souza. On Conflict-Free Spanning Tree: Mapping tractable and hard instances through the lenses of graph classes. Theoretical Computer Science, 1031(??):??, March 21, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000192.

Vaezi:2025:VER

[122] Arash Vaezi, Bodhayan Roy, and Mohammad Ghodsi. Visibility extension via reflection. Theoretical Computer Science, 1031(??):??, March 21, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000210.

Dutta:2025:SDC

[123] Subhranil Dutta, Ratna Dutta, and Sourav Mukhopadhyay. Securing data in the cloud using pairing-free inner product functional encryption with unbounded vector size. Theoretical Computer Science, 1031(??): ??, March 21, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000234.

Anonymous:2025:EBi

[124] Anonymous. Editorial Board. Theoretical Computer Science, 1031(??):

??, March 21, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000544.

Anonymous:2025:Mc

[125] Anonymous. 21 March 2025. Theoretical Computer Science, 1031(??):??, March 21, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cohen:2025:PLR

[126] Keren Cohen. Powers of low rank sparse matrices. Theoretical Computer Science, 1032(??):??, March 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000209.

Gale:2025:CMO

[127] Ella Gale, Leo Lobski, and Fabio Zanasi. A categorical model for organic chemistry. Theoretical Computer Science, 1032(??):??, March 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000222.

Franklin:2025:CCC

[128] Johanna N. Y. Franklin, Rupert Hölzl, Alexander Melnikov, Keng Meng Ng, and Daniel Turetsky. Computable classifications of continuous, transducer, and regular functions. Theoretical Computer Science, 1032(??): ??, March 29, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000246.

Gadouleau:2025:LPC

[129] Maximilien Gadouleau, George B.

Mertzios, and Viktor Zamaraev. Linear Programming complementation. Theoretical Computer Science, 1032 (??):??, March 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000258.

Brettell:2025:CSV

[130] Nick Brettell, Jelle J. Oostveen, Sukanya Pandey, Daniël Paulusma, Johannes Rauch, and Erik Jan van Leeuwen. Computing subset vertex covers in Hfree graphs. Theoretical Computer Science, 1032(??):??, March 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500026X.

Cyprys:2025:SMH

[131] Paweł Cyprys, Shlomi Dolev, and Shlomo Moran. Self-masking for hardening inversions. Theoretical Computer Science, 1032(??):??, March 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0304397525000325.

Lu:2025:DPG

[132] Zenan Lu, Wanyue Xu, and Zhongzhi Zhang. Diagonal of pseudoinverse of graph Laplacian: Fast estimation and exact results. Theoretical Computer Science, 1032(??):??, March 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000404.

Brandstadt:2025:ISM

[133] Andreas Brandstädt, Vadim Lozin, and Raffaele Mosca. Independent sets of maximum weight beyond clawfree graphs and related problems. Theoretical Computer Science, 1032 (??):??, March 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000416.

Yin:2025:USS

[134] Bai Yin, Qianru Zhou, Hai Liu, Yan Wang, Baolei Cheng, and Jianxi Fan. Unpaired set-to-set disjoint path routings in recursive match networks. Theoretical Computer Science, 1032 (??):??, March 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000490.

Anonymous:2025:EBj

[135] Anonymous. Editorial Board. Theoretical Computer Science, 1032(??): ??, March 29, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000696.

Anonymous:2025:Md

[136] Anonymous. 29 March 2025. Theoretical Computer Science, 1032(??):??, March 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Madireddy:2025:GRB

[137] Raghunath Reddy Madireddy, Subhas C. Nandy, and Supantha Pandit.

On the geometric red–blue set cover problem. Theoretical Computer Science, 1033(??):??, April 7, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000271.

Zhu:2025:OMS

[138] Liwang Zhu and Zhongzhi Zhang.

Opinion maximization in social networks via link recommendation. Theoretical Computer Science, 1033(??): ??, April 7, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000283.

Bousquet:2025:SCS

[139] Nicolas Bousquet and Sébastien Zeitoun.

A subquadratic certification scheme for P_{5}-free graphs. Theoretical Computer Science, 1033(??):??, April 7, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000295.

Hanaka:2025:FMS

[140] Tesshu Hanaka and Yasuaki Kobayashi.

Finding a minimum spanning tree with a small non-terminal set. Theoretical Computer Science, 1033(??): ??, April 7, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000301.

Hara:2025:LRS

[141] Keisuke Hara. A linkable ring signature scheme with unconditional anonymity in the standard model. Theoretical Computer Science, 1033 (??):??, April 7, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000313.

Santo:2025:HAC

[142] José Espírito Santo and Gilda Ferreira. How to avoid the commuting conversions of IPC. Theoretical Computer Science, 1033(??):??, April 7, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000398.

Estermann:2025:ORA

[143] Benjamin Estermann, Stefan Kramer, Roger Wattenhofer, and Kanye Ye Wang. Optimizing resource allocation: an active learning approach to iterative combinatorial auctions. Theoretical Computer Science, 1033 (??):??, April 7, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000428.

Anonymous:2025:EBk

[144] Anonymous. Editorial Board. Theoretical Computer Science, 1033(??): ??, April 7, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000751.

Anonymous:2025:Aa

[145] Anonymous. 7 April 2025. Theoretical Computer Science, 1033(??):??, April 7, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2025:NQD

[146] Qisheng Wang. A note on quantum divide and conquer for minimal string rotation. Theoretical Computer Science, 1034(??):??, April 22, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000581.

Michel:2025:LBG

[147] Lukas Michel and Alex Scott. Lower bounds for graph reconstruction with maximal independent set queries. Theoretical Computer Science, 1034(??): ??, April 22, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000593.

Hu:2025:SSE

[148] Shengrong Hu. Shell sort with expected complexity of O(n log_{2} n). Theoretical Computer Science, 1034 (??):??, April 22, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500060X.

Liang:2025:NCG

[149] Zihui Liang, Bakh Khoussainov, and Mingyu Xiao. Network control games played on graphs. Theoretical Computer Science, 1034(??): ??, April 22, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000611.

Bergstra:2025:RNS

[150] Jan A. Bergstra and John V. Tucker. For rational numbers with Suppes– Ono division, equational validity is oneone equivalent with Diophantine unsolvability. Theoretical Computer Science, 1034(??):??, April 22, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000623.

Dhawan:2025:FSV

[151] Abhishek Dhawan. Fast and simple (1 + ε)Δ-edge-coloring of dense graphs. Theoretical Computer Science, 1034(??):??, April 22, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000647.

Anonymous:2025:EBl

[152] Anonymous. Editorial Board. Theoretical Computer Science, 1034(??): ??, April 22, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525001033.

Anonymous:2025:Ab

[153] Anonymous. 22 April 2025. Theoretical Computer Science, 1034(??):??, April 22, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bhagat:2025:MMM

[154] Subhash Bhagat, Krishnendu Mukhopadhyaya, and Rajarshi Ray. The min-move mutual visibility problem for disoriented asynchronous robots. Theoretical Computer Science, 1035 (??):??, April 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000635.

Shaw:2025:PQS

[155] Surbhi Shaw and Ratna Dutta. Postquantum secure compact deterministic wallets from isogeny-based signatures with rerandomized keys. Theoretical Computer Science, 1035(??): ??, April 29, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000659.

Borowiecki:2025:DEG

[156] Piotr Borowiecki, Shantanu Das, Dariusz Dereniowski, and Łukasz Kuszner. Discrete evacuation in graphs with multiple exits. Theoretical Computer Science, 1035(??):??, April 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000799.

DiCicco:2025:ISB

[157] Mason DiCicco, Henry Poskanzer, and Daniel Reichman. Inoculation strategies for bounded degree graphs. Theoretical Computer Science, 1035 (??):??, April 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000805.

Anonymous:2025:EBm

[158] Anonymous. Editorial Board. Theoretical Computer Science, 1035(??): ??, April 29, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525001094.

Anonymous:2025:Ac

[159] Anonymous. 29 April 2025. Theoretical Computer Science, 1035(??):??, April 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2025:SDB

[160] Zizhen Zhang, Deying Li, Yongcai Wang, Wenping Chen, and Yuqing Zhu. Sequential decision based learning method for influence maximization. Theoretical Computer Science, 1036(??):??, May 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000854.

Jia:2025:NSC

[161] Zhenhua Jia and Qingguo Li. Notes on Smyth-completes and local Yonedacompletes. Theoretical Computer Science, 1036(??):??, May 3, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000866.

Gomes:2025:WCM

[162] Guilherme C. M. Gomes, Bruno P. Masquio, Paulo E. D. Pinto, Vinicius F. dos Santos, and Jayme L. Szwarcfiter. Weighted connected matchings. Theoretical Computer Science, 1036(??):??, May 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000878.

Tao:2025:BER

[163] Liangde Tao, Lin Chen, Lei Xu, Weidong Shi, Md Mahabub Uz Zaman, and Ahmed Sunny. Bribery in elections with randomly selected voters: Hardness and algorithm. Theoretical Computer Science, 1036(??):??, May 3, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500088X.

Manurangsi:2025:OMG

[164] Pasin Manurangsi and Warut Suksompong. Ordinal maximin guarantees for group fair division. Theoretical Computer Science, 1036(??): ??, May 3, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000891.

Feletti:2025:CPA

[165] Caterina Feletti, Lucia Mambretti, Carlo Mereghetti, and Beatrice Palano. Computational power of autonomous robots: Transparency vs. opaqueness. Theoretical Computer Science, 1036(??):??, May 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500091X.

Amrane:2025:CDP

[166] Amazigh Amrane, Hugo Bazille, Uli Fahrenberg, and Krzysztof Ziemiański. Closure and decision properties for higher-dimensional automata. Theoretical Computer Science, 1036(??): ??, May 3, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000945.

Anonymous:2025:EBn

[167] Anonymous. Editorial Board. Theoretical Computer Science, 1036(??): ??, May 3, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525001343.

Anonymous:2025:Me

[168] Anonymous. 3 May 2025. Theoretical Computer Science, 1036(??):??, May 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bose:2025:ESR

[169] Prosenjit Bose, Jean-Lou De Carufel, and Sandrine Njoo. The exact spanning ratio of the parallelogram Delaunay graph. Theoretical Computer Science, 1037(??):??, May 11, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525000908.

Hellmuth:2025:SNH

[170] Marc Hellmuth and Guillaume E. Scholz. Solving NP-hard problems on GaTEx graphs: Linear-time algorithms for perfect orderings, cliques, colorings, and independent sets. Theoretical Computer Science, 1037(??): ??, May 11, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000957.

Lingas:2025:MMM

[171] Andrzej Lingas and Mia Persson.

(min, +) matrix and vector products for inputs decomposable into few monotone subsequences. Theoretical Computer Science, 1037(??):??, May 11, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000969.

Koiran:2025:CDS

[172] Pascal Koiran and Subhayan Saha.

Complete decomposition of symmetric tensors in linear time and polylogarithmic precision. Theoretical Computer Science, 1037(??):??, May 11, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000970.

Bose:2025:PGB

[173] Prosenjit Bose, Jean-Lou De Carufel, Anil Maheshwari, and Karthik Murali. On 1-planar graphs with bounded copnumber. Theoretical Computer Science, 1037(??):??, May 11, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000982.

Petrovic:2025:KLB

[174] Tomislav Petrović. Kolmogorov– Loveland betting strategies lose the betting game on open sets. Theoretical Computer Science, 1037(??): ??, May 11, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752500115X.

Niu:2025:SNR

[175] Bei Niu. Some new results on equitable tree-coloring parameters of graphs. Theoretical Computer Science, 1037 (??):??, May 11, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525001161.

Anonymous:2025:EBo

[176] Anonymous. Editorial Board. Theoretical Computer Science, 1037(??): ??, May 11, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525001446.

Anonymous:2025:Mf

[177] Anonymous. 11 May 2025. Theoretical Computer Science, 1037(??):??, May 11, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutner:2025:RRD

[178] David C. Kutner and Iain A. Stewart. Reconfigurable routing in data center networks. Theoretical Computer Science, 1038(??):??, May 22, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000921.

Maire:2025:CZK

[179] Jules Maire and Damien Vergnaud.

Compact zero-knowledge arguments for Blum integers. Theoretical Computer Science, 1038(??):??, May 22, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000933.

Rosenkrantz:2025:TFP

[180] Daniel J. Rosenkrantz, Madhav V.

Marathe, Zirou Qiu, and S. S. Ravi. Theoretical foundations for parent divorcing transformations in Bayesian networks. Theoretical Computer Science, 1038(??):??, May 22, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001148.

Sharma:2025:MOA

[181] Avisek Sharma, Satakshi Ghosh, Pritam Goswami, and Buddhadeb Sau. Move-optimal arbitrary pattern formation by mobile robots on rectangular grid using near-optimal spatial area. Theoretical Computer Science, 1038(??):??, May 22, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001173.

Ng:2025:SSC

[182] Chii Liang Ng, Denis C. K. Wong, Gek L. Chia, Bok Min Goi, Wai Kong Lee, and Wun She Yap. A signature scheme constructed from zero knowledge argument of knowledge for the subgraph isomorphism problem. Theoretical Computer Science, 1038 (??):??, May 22, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001185.

Stramaglia:2025:USV

[183] Anna Stramaglia, Jeroen J. A. Keiren, and Thomas Neele. Unfolding state variables improves model checking performance. Theoretical Computer Science, 1038(??):??, May 22, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001197.

Behrouz:2025:BFT

[184] Pourandokht Behrouz, Orestis Konstantinidis, Nikos Leonardos, Aris Pagourtzis, Ioannis Papaioannou, and Marianna Spyrakou. Byzantine faulttolerant protocols for (n, f)-evacuation from a circle. Theoretical Computer Science, 1038(??):??, May 22, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001239.

Anonymous:2025:EBp

[185] Anonymous. Editorial Board. Theoretical Computer Science, 1038(??): ??, May 22, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001628.

Anonymous:2025:Mg

[186] Anonymous. 22 May 2025. Theoretical Computer Science, 1038(??):??, May 22, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2025:TAS

[187] Anonymous. Tribute to Arto Salomaa:

a towering figure in theoretical computer science. Theoretical Computer Science, 1039(??):??, June 4, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525000507.

Coudert:2025:SSP

[188] David Coudert, Andrea D’Ascenzo, and Clément Rambaud. k-shortest simple paths in bounded treewidth graphs. Theoretical Computer Science, 1039(??):??, June 4, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001203.

Duvignau:2025:SSM

[189] Romaric Duvignau, Michel Raynal, and Elad Michael Schiller. Self-stabilizing multivalued consensus in the presence of Byzantine faults and asynchrony. Theoretical Computer Science, 1039(??):??, June 4, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001227.

Rosner:2025:CSG

[190] Shaul Rosner and Tami Tamir. Costsharing games with rank-based utilities. Theoretical Computer Science, 1039(??):??, June 4, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001240.

Zehmakan:2025:ODG

[191] Ahad N. Zehmakan. Opinion diffusion in graphs: an adversarial approach. Theoretical Computer Science, 1039(??):??, June 4, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001264.

Chen:2025:RSN

[192] Qigong Chen, Jiafei Liu, Chia-Wei Lee, Jingli Wu, and Gaoshi Li. The reliability of (n, k)-star network in terms of non-inclusive fault pattern. Theoretical Computer Science, 1039 (??):??, June 4, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001276.

Bhattacharya:2025:IAO

[193] Binay Bhattacharya, Mordecai J. Golin, Yuya Higashikawa, Tsunehiko Kameda, and Naoki Katoh. Improved algorithms for optimal k sink location on path networks. Theoretical Computer Science, 1039(??):??, June 4, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001288.

Halle:2025:TFG

[194] Sylvain Hallé and Hugo Tremblay. A theory of fine-grained lineage for functions on structured objects. Theoretical Computer Science, 1039(??): ??, June 4, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001306.

Anonymous:2025:EBq

[195] Anonymous. Editorial Board. Theoretical Computer Science, 1039(??): ??, June 4, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500180X.

Anonymous:2025:Jc

[196] Anonymous. 4 June 2025. Theoretical Computer Science, 1039(??):??, June 4, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Georgiou:2025:TSI

[197] Konstantinos Georgiou and Evangelos Kranakis. TCS special issue on selected papers from AlgoWin 2023. Theoretical Computer Science, 1040 (??):??, June 22, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001215.

Furer:2025:EDS

[198] Martin Fürer, Carlos Hoppen, and Vilmar Trevisan. Efficient diagonalization of symmetric matrices associated with graphs of small treewidth. Theoretical Computer Science, 1040 (??):??, June 22, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001252.

vanHeerdt:2025:CLS

[199] Gerco van Heerdt, Justin Hsu, Joël Ouaknine, and Alexandra Silva. Convex language semantics for nondeterministic probabilistic automata. Theoretical Computer Science, 1040(??): ??, June 22, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500129X.

Koponen:2025:RET

[200] Vera Koponen and Yasmin Tousinejad. Random expansions of trees with bounded height. Theoretical Computer Science, 1040(??):??, June 22, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001392.

Fang:2025:IVM

[201] Nan Fang, Lu Liu, and Liang Yu. Integer-valued martingales and cl-Turing reductions. Theoretical Computer Science, 1040(??):??, June 22, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0304397525001409.

Lohani:2025:ECC

[202] Rohit Lohani and Krishnendra Shekhawat. Existence and construction of a Cshaped module within a floorplan. Theoretical Computer Science, 1040 (??):??, June 22, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001471.

Guo:2025:ELD

[203] Huiling Guo, Shurong Zhang, Shanshan Shan, Lin Chen, and Weihua Yang. The extra local diagnosability and diagnosis algorithm of networks under the PMC model. Theoretical Computer Science, 1040(??):??, June 22, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001483.

Anonymous:2025:EBr

[204] Anonymous. Editorial Board. Theoretical Computer Science, 1040(??): ??, June 22, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001926.

Anonymous:2025:Jd

[205] Anonymous. 22 June 2025. Theoretical Computer Science, 1040(??):??, June 22, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2025:PBA

[206] Houshuang Chen, Yuchen He, and Chihao Zhang. On the problem of Best Arm Retention. Theoretical Computer Science, 1041(??):??, July 7, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001513.

Bhargav:2025:LBS

[207] C. S. Bhargav, Prateek Dwivedi, and Nitin Saxena. Lower bounds for the sum of small-size algebraic branching programs. Theoretical Computer Science, 1041(??):??, July 7, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001525.

Liu:2025:IKP

[208] Yuxi Liu and Mingyu Xiao. An improved kernel and parameterized algorithm for deletion to induced matching. Theoretical Computer Science, 1041(??):??, July 7, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001537.

Charalampopoulos:2025:SCW

[209] Panagiotis Charalampopoulos, Solon P. Pissis, Jakub Radoszewski, Wojciech Rytter, Tomasz Waleń, and Wiktor Zuba. Subsequence covers of words. Theoretical Computer Science, 1041(??):??, July 7, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001549.

Capobianco:2025:JGE

[210] Silvio Capobianco. Not just the Garden of Eden: Additional takes on cellular automata surjectivity. Theoretical Computer Science, 1041(??): ??, July 7, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001550.

Chen:2025:EPM

[211] Meirun Chen, Cheng-Kuan Lin, and Stafen Wang. An efficient PMC modelbased local diagnosis structure and algorithm. Theoretical Computer Science, 1041(??):??, July 7, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001562.

Perrotin:2025:FSP

[212] Pacôme Perrotin, Eurico Ruivo, and Pedro Paulo Balbi. Fast solutions to k-parity and k-synchronisation using parallel automata networks. Theoretical Computer Science, 1041(??): ??, July 7, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001574.

Dubois:2025:WSY

[213] Swan Dubois, Laurent Feuilloley, Franck Petit, and Mikaël Rabie. When should you wait before updating? — Toward a robustness refinement. Theoretical Computer Science, 1041(??): ??, July 7, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001586.

Tani:2025:QAF

[214] Seiichiro Tani. Quantum algorithm for finding the optimal variable ordering for binary decision diagrams. Theoretical Computer Science, 1041 (??):??, July 7, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525001689.

Anonymous:2025:EBs

[215] Anonymous. Editorial Board. Theoretical Computer Science, 1041(??): ??, July 7, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002245.

Anonymous:2025:Je

[216] Anonymous. 7 July 2025. Theoretical Computer Science, 1041(??):??, July 7, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ceccherini-Silberstein:2025:SFM

[217] Tullio Ceccherini-Silberstein, Michel Coornaert, and Xuan Kien Phung. Stable finiteness of monoid algebras and surjunctivity. Theoretical Computer Science, 1042(??):??, July 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525001665.

Goles:2025:DST

[218] Eric Goles, Pedro Montealegre, Martín Ríos-Wilson, and Sylvain Sené. Dynamical stability of threshold networks over undirected signed graphs. Theoretical Computer Science, 1042 (??):??, July 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525001677.

Saez:2025:DFG

[219] Pablo Sáez. Deciding the feasibility of a given throughput rate in 2-cyclic robotic scheduling. Theoretical Computer Science, 1042(??):??, July 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001719.

Beaton:2025:VCG

[220] Iain Beaton and Ben Cameron. Vertexcritical graphs in co-gem-free graphs. Theoretical Computer Science, 1042 (??):??, July 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001720.

Bhojraj:2025:NEQ

[221] Tejas Bhojraj. von Neumann entropy and quantum algorithmic randomness. Theoretical Computer Science, 1042(??):??, July 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001732.

Miao:2025:MCS

[222] Qihui Miao, Yongshun Shen, Wengke Zang, and Yuzhen Zhao. Multiple channels spiking neural P systems with lateral inhibition, rules dynamic generation and removal. Theoretical Computer Science, 1042(??):??, July 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001744.

Salo:2025:SCP

[223] Ville Salo and Ilkka Törmä. Structure and computability of preimages in the Game of Life. Theoretical Computer Science, 1042(??):??, July 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525001756.

Tao:2025:EEP

[224] Biaoshuai Tao, Xiaowei Wu, Ziqi Yu, and Shengwei Zhou. On the existence of EFX (and Pareto-optimal) allocations for binary chores. Theoretical Computer Science, 1042(??):??, July 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525001860.

Calamoneri:2025:EVC

[225] Tiziana Calamoneri and Federico Corò. (Eternal) vertex cover numbers of infinite and finite grid graphs. Theoretical Computer Science, 1042(??): ??, July 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525001999.

Anonymous:2025:EBt

[226] Anonymous. Editorial Board. Theoretical Computer Science, 1042(??): ??, July 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002348.

Anonymous:2025:Jf

[227] Anonymous. 19 July 2025. Theoretical Computer Science, 1042(??):??, July 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pombo:2025:DTL

[228] Carlos G. Lopez Pombo, Agustín E.

Martinez-Suñé, and Emilio Tuosto. A dynamic temporal logic for quality of service in choreographic models. Theoretical Computer Science, 1043(??):??, July 30, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001859.

Chen:2025:IAS

[229] Haiming Chen, Hong Huang, Rongchen Li, Chengyao Peng, and Weihao Su. Incremental algorithms for solving regular expression intersection non-emptiness. Theoretical Computer Science, 1043 (??):??, July 30, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001872.

Karakostas:2025:AAM

[230] George Karakostas and Stavros G. Kolliopoulos. Approximation algorithms for maximum weighted throughput on unrelated machines. Theoretical Computer Science, 1043(??):??, July 30, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001963.

Navarro:2025:RMB

[231] Gonzalo Navarro and Cristian Urbina. Repetitiveness measures based on string morphisms. Theoretical Computer Science, 1043(??):??, July 30, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001975.

Bhyravarapu:2025:DCR

[232] Sriram Bhyravarapu, Swati Kumari, and I. Vinod Reddy. Dynamic coloring on restricted graph classes. Theoretical Computer Science, 1043 (??):??, July 30, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525001987.

Bajaj:2025:CPD

[233] Shivam Bajaj, Eric Torng, and Shaunak D. Bopardikar. Competitive perimeter defense in linear environments. Theoretical Computer Science, 1043(??):??, July 30, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002002.

Zhao:2025:MCV

[234] Jingyang Zhao and Mingyu Xiao. Multidepot capacitated vehicle routing with improved approximation guarantees. Theoretical Computer Science, 1043(??):??, July 30, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002038.

Anonymous:2025:EBu

[235] Anonymous. Editorial Board. Theoretical Computer Science, 1043(??): ??, July 30, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002440.

Anonymous:2025:Jg

[236] Anonymous. 30 July 2025. Theoretical Computer Science, 1043(??):??, July 30, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Voogd:2025:CSE

[237] Erik Voogd, Åsmund Aqissiaq Arild Kløvstad, Einar Broch Johnsen, and Andrzej Wa̧sowski. Compositional symbolic execution semantics. Theoretical Computer Science, 1044 (??):??, August 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002014.

Martinez:2025:MMA

[238] Conrado Martínez, Cyril Nicaud, and Pablo Rotondo. Mathematical models to analyze Lua hybrid tables. Theoretical Computer Science, 1044 (??):??, August 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002026.

Guo:2025:TED

[239] Xinru Guo, Sijia Dai, Xinxin Han, Yicheng Xu, and Yong Zhang. Time and energy driven online scheduling problem in EV charging. Theoretical Computer Science, 1044(??): ??, August 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500204X.

Araujo:2025:ACG

[240] Samuel N. Araújo, João Marcos Brito, Raquel Folz, Rosiane de Freitas, and Rudini M. Sampaio. Algorithms and complexity of graph convexity partizan games. Theoretical Computer Science, 1044(??):??, August 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002051.

Zhang:2025:ARR

[241] Shasha Zhang, Jie Zhou, Yanfang Ma, and Haiyu Pan. Alternating refinement relations for fuzzy concurrent game structures. Theoretical Computer Science, 1044(??):??, August 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002063.

Lee:2025:MLC

[242] Chien-Ting Lee, Chang-Biau Yang, and Kuo-Si Huang. The merged longest common increasing subsequence problem. Theoretical Computer Science, 1044(??):??, August 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002075.

Abhinav:2025:TTF

[243] Ankit Abhinav, Satyabrata Jana, and Abhishek Sahu. Towards transitive-free digraphs. Theoretical Computer Science, 1044(??):??, August 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002087.

Amiri:2025:CCA

[244] Saeed Akhoondian Amiri, Alexandru Popa, Mohammad Roghani, Golnoosh Shahkarami, Reza Soltani, and Hossein Vahidi. Complexity of computing the anti-Ramsey numbers for paths. Theoretical Computer Science, 1044 (??):??, August 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002099.

Kshemkalyani:2025:DMR

[245] Ajay D. Kshemkalyani. Dispersion of mobile robots on graphs in the asynchronous model. Theoretical Computer Science, 1044(??): ??, August 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002105.

Shi:2025:RPE

[246] Feng Shi, Daoyu Huang, Xiankun Yan, and Frank Neumann. Runtime performance of evolutionary algorithms for the chance-constrained makespan scheduling problem. Theoretical Computer Science, 1044(??):??, August 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002129.

Lv:2025:MCS

[247] Xiaowei Lv, Yongcai Wang, and Deying Li. Maximum core spanning tree maintenance for large dynamic graphs. Theoretical Computer Science, 1044 (??):??, August 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002166.

Anonymous:2025:EBv

[248] Anonymous. Editorial Board. Theoretical Computer Science, 1044(??): ??, August 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002543.

Anonymous:2025:Ad

[249] Anonymous. 1 August 2025. Theoretical Computer Science, 1044(??):??, August 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ikebuchi:2025:CPS

[250] Mirai Ikebuchi. Cyclic proofs and size-change termination. Theoretical Computer Science, 1045(??):??, August 11, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002117.

Liu:2025:FFE

[251] Zhe Liu, Wenguo Yang, and Suixiang Gao. Finding fair and efficient allocations of indivisible chores. Theoretical Computer Science, 1045 (??):??, August 11, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002130.

Chen:2025:UBS

[252] Boxi Chen, Jingfang Su, and Hongwei Du. UAV-based sweep coverage for timesensitive targets with restricted visible areas. Theoretical Computer Science, 1045(??):??, August 11, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002142.

Li:2025:TGA

[253] Yifei Li, Hao Ye, and Hejiao Huang.

Task group allocation for multi-load agent pickup and delivery problem. Theoretical Computer Science, 1045 (??):??, August 11, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002154.

Junnila:2025:LSR

[254] Ville Junnila, Tero Laihonen, and Tuomo Lehtilä. Levenshtein’s sequence reconstruction problem and results for larger alphabet sizes. Theoretical Computer Science, 1045(??): ??, August 11, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002178.

Alajaji:2025:MUW

[255] John Alajaji and Kai Salomaa. Maximal universal width of an AFA is NPhard. Theoretical Computer Science, 1045(??):??, August 11, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752500218X.

Ge:2025:NFD

[256] Gaotao Ge, Jiafei Liu, Dajin Wang, Jingli Wu, and Gaoshi Li. A novel fault diagnostic algorithm with multiple characteristics for multiprocessor systems. Theoretical Computer Science, 1045(??):??, August 11, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002191.

Tale:2025:TBG

[257] Prafullkumar Tale. Telephone broadcast on graphs of treewidth two. Theoretical Computer Science, 1045 (??):??, August 11, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002208.

Bruckner:2025:PCL

[258] Guido Brückner and Ignaz Rutter. Partial and constrained level planarity. Theoretical Computer Science, 1045 (??):??, August 11, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002294.

Jahier:2025:MCD

[259] Erwan Jahier, Karine Altisen, Stéphane Devismes, and Gabriel B. Sant’Anna. Model checking of distributed algorithms using synchronous programs. Theoretical Computer Science, 1045 (??):??, August 11, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002300.

Ascone:2025:CGA

[260] Rocco Ascone, Giulia Bernardini, and Luca Manzoni. Cycles and global attractors of reactantless and inhibitorless reaction systems. Theoretical Computer Science, 1045(??): ??, August 11, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002385.

Anonymous:2025:EBw

[261] Anonymous. Editorial Board. Theoretical Computer Science, 1045(??): ??, August 11, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002786.

Anonymous:2025:Ae

[262] Anonymous. 11 August 2025. Theoretical Computer Science, 1045(??):??, August 11, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daripa:2025:HPP

[263] Krishna Daripa and Alak Kumar Datta. Hamiltonian path in permutation graphs. Theoretical Computer Science, 1046(??):??, August 28, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002397.

Wang:2025:AAM

[264] Yuanyuan Wang, Qingqin Nong, Wenjing Liu, and Qizhi Fang. Approximation algorithm and mechanism design for bisubmodular welfare maximization problem. Theoretical Computer Science, 1046(??):??, August 28, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002403.

Barenboim:2025:SOE

[265] Leonid Barenboim and Tzalik Maimon. Sampling and output estimation in distributed algorithms and LCAs. Theoretical Computer Science, 1046 (??):??, August 28, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500249X.

Wan:2025:EDA

[266] Zongqi Wan, Jialin Zhang, Xiaoming Sun, and Zhijie Zhang. Efficient deterministic algorithms for maximizing symmetric submodular functions. Theoretical Computer Science, 1046(??):??, August 28, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002506.

Kourtis:2025:MFP

[267] Georgios Kourtis, Clare Dixon, and Michael Fisher. Monodic fragments of probabilistic first-order temporal logic with bounded semantics. Theoretical Computer Science, 1046(??): ??, August 28, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002579.

Fleischmann:2025:GNW

[268] Pamela Fleischmann, Annika Huch, and Dirk Nowotka. Generalised Nyldon words. Theoretical Computer Science, 1046(??):??, August 28, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002580.

Cardi:2025:WCF

[269] Pierre Cardi, Laurent Gourvès, and Julien Lesca. Worst-case fair guarantees when spending a common budget. Theoretical Computer Science, 1046(??):??, August 28, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002609.

Brosse:2025:CSF

[270] Caroline Brosse, Nicolas Martins, Nicolas Nisse, and Rudini Sampaio. The Convex Set Forming Game. Theoretical Computer Science, 1046(??): ??, August 28, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002610.

Anonymous:2025:EBx

[271] Anonymous. Editorial Board. Theoretical Computer Science, 1046(??): ??, August 28, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002907.

Anonymous:2025:Af

[272] Anonymous. 28 August 2025. Theoretical Computer Science, 1046(??):??, August 28, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2025:LTF

[273] Weian Li and Yu Zhou. Locating two facilities on a square with a minimum distance requirement. Theoretical Computer Science, 1047(??): ??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002488.

Arvind:2025:ERC

[274] V. Arvind, Johannes Köbler, and Oleg Verbitsky. On the expressibility of the reconstructional color refinement. Theoretical Computer Science, 1047 (??):??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002592.

Khoo:2025:DLO

[275] Kai Jun Khoo, Heer Tern Koh, and Keng Meng Ng. A discrete linear order with non-dense punctual degrees.

Theoretical Computer Science, 1047 (??):??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002622.

Aceto:2025:AWB

[276] Luca Aceto, Valentina Castiglioni, Anna Ingólfsdóttir, and Bas Luttik. Axiomatising weak bisimulation congruences over CCS with left merge and communication merge. Theoretical Computer Science, 1047(??): ??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002634.

Alpay:2025:SMM

[277] Daniel Alpay and Liora Mayats-Alpay.

Similarity metrics, metrics, and conditionally negative definite functions. Theoretical Computer Science, 1047 (??):??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002646.

Romero:2025:CLM

[278] Pablo Romero. Characterization of locally most split reliable graphs. Theoretical Computer Science, 1047(??): ??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002658.

Souza:2025:DTS

[279] Victor F. C. Souza, Ferdinando Cicalese, Eduardo Sany Laber, and Marco Molinaro. Decision trees with short explainable rules. Theoretical Computer Science, 1047(??): ??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002828.

Tan:2025:LCH

[280] Xuehou Tan and Rong Chen. Largest convex hulls for convex-hull disjoint clusters with bounded size. Theoretical Computer Science, 1047(??): ??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752500283X.

Anonymous:2025:EPJ

[281] Anonymous. Editorial: a personal journey with TCS. Theoretical Computer Science, 1047(??): ??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002853.

Anonymous:2025:TBE

[282] Anonymous. TCS B editorial. Theoretical Computer Science, 1047(??): ??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525002865.

Anonymous:2025:EBy

[283] Anonymous. Editorial Board. Theoretical Computer Science, 1047(??): ??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003147.

Anonymous:2025:Sa

[284] Anonymous. 3 September 2025. Theoretical Computer Science, 1047(??):??, September 3, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asahiro:2025:CCA

[285] Yuichi Asahiro and Masafumi Yamashita. Compatibility of convergence algorithms for autonomous mobile robots. Theoretical Computer Science, 1048(??):??, September 13, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500266X.

Civril:2025:RNN

[286] A. Çivril. Retraction notice to “A New Approximation Algorithm for the Minimum 2-Edge-Connected Spanning Subgraph Problem” [Theoretical Computer Science 943 (2023) 121–130]. Theoretical Computer Science, 1048(??): ??, September 13, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002841. See

[2].

Wei:2025:LEU

[287] Qi Wei, Xuehou Tan, Haonan Wu, Xiaolin Yao, and Yonggong Ren. Online exploration of an unbounded region with one obstacle. Theoretical Computer Science, 1048(??): ??, September 13, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500297X.

Hirota:2025:EAE

[288] Miyuji Hirota and Yoshifumi Sakai. Efficient algorithms for enumerating maximal common subsequences of two strings. Theoretical Computer Science, 1048(??):??, September 13, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002981.

Couto:2025:MDB

[289] Fernanda Couto, Luís Cunha, Edmundo Pinto, and Daniel Posner. Minimizing distances between vertices and edges through tree t-spanners. Theoretical Computer Science, 1048(??): ??, September 13, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003007.

Hori:2025:GCG

[290] Sawaka Hori and Kenji Mikawa. Gray code generation of binary trees encoded in left-child sequences. Theoretical Computer Science, 1048(??): ??, September 13, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003044.

Anonymous:2025:EBz

[291] Anonymous. Editorial Board. Theoretical Computer Science, 1048(??): ??, September 13, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003251.

Anonymous:2025:Sb

[292] Anonymous. 13 September 2025. Theoretical Computer Science, 1048(??):??, September 13, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ashok:2025:BW

[293] Pradeesha Ashok, Sayani Das, Lawqueen Kanesh, Saket Saurabh, Avi Tomar, and Shaily Verma. Burn and win. Theoretical Computer Science, 1049(??): ??, September 17, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002956.

Kammer:2025:SRS

[294] Frank Kammer, Johannes Meintrup, and Andrej Sajenko. Sorting and ranking of self-delimiting numbers with applications to outerplanar graph isomorphism. Theoretical Computer Science, 1049(??):??, September 17, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002968.

D:2025:FPR

[295] Karthika D., Muthucumaraswamy R., Sriram Bhyravarapu, Satyabrata Jana, and Saket Saurabh. Further parameterized results on weak Grundy coloring. Theoretical Computer Science, 1049(??):??, September 17, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525002993.

Verrill:2025:SBP

[296] H. A. Verrill. L-systems for the boundaries of plane-filling folding curves. Theoretical Computer Science, 1049(??): ??, September 17, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003019.

Mitchell:2025:SGO

[297] Joseph S. B. Mitchell and Supantha Pandit. On some geometric optimization problems with segments. Theoretical Computer Science, 1049(??): ??, September 17, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003020.

Chakraborty:2025:AAA

[298] Dibyayan Chakraborty and Yann Vaxès. Additive approximation algorithm for geodesic centers in δ-hyperbolic graphs. Theoretical Computer Science, 1049 (??):??, September 17, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003032.

Duan:2025:SBB

[299] Zhenhua Duan, Cong Tian, Nan Zhang, Chaofeng Yu, Mengfei Yang, and Jia He. SAT-based bounded model checking for propositional projection temporal logic. Theoretical Computer Science, 1049(??):??, September 17, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003068.

Xu:2025:AAM

[300] Yuxuan Xu, Yingli Ran, and Zhao Zhang. A 1/2-approximation algorithm for maximum interval multicover. Theoretical Computer Science, 1049(??):??, September 17, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500307X.

Yuan:2025:SPB

[301] Jing Yuan and Shaojie Tang. Submodular participatory budgeting. Theoretical Computer Science, 1049(??): ??, September 17, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003081.

Liu:2025:MSV

[302] Jingyi Liu, Yixin Cao, Ling Gai, and Jianxin Wang. Minimum sum vertex cover: Difficulty of ordering. Theoretical Computer Science, 1049(??): ??, September 17, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003093.

Cao:2025:ECS

[303] Tianyue Cao, Zhao Jin, Bowen Zhang, Yongzhi Cao, and Hanpin Wang. Expressive completeness of separation logic in block-based cloud storage systems. Theoretical Computer Science, 1049 (??):??, September 17, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752500310X.

Mizera:2025:PSD

[304] Andrzej Mizera and Jakub Zarzycki. pbn-STAC: Deep reinforcement learning-based framework for cellular reprogramming. Theoretical Computer Science, 1049(??):??, September 17, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003202.

Anonymous:2025:EBba

[305] Anonymous. Editorial Board. Theoretical Computer Science, 1049(??): ??, September 17, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003810.

Anonymous:2025:Sc

[306] Anonymous. 17 September 2025. Theoretical Computer Science, 1049(??):??, September 17, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Saxena:2025:PCD

[307] Ashish Saxena and Kaushik Mondal. Path connected dynamic graphs with a study of dispersion and exploration. Theoretical Computer Science, 1050 (??):??, September 27, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003287.

Joshi:2025:MAP

[308] G. Joshi and D. Rust. Monochromatic arithmetic progressions in the Fibonacci, Thue–Morse, and Rudin–Shapiro words. Theoretical Computer Science, 1050 (??):??, September 27, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003299.

Wang:2025:FDF

[309] Ruida Wang, Xianhui Lu, Yundi Wen, Zhihao Li, Benqiang Wei, Kunpeng Wang, and Lixia Luo. Full domain functional bootstrapping using the prime cyclotomic ring. Theoretical Computer Science, 1050(??): ??, September 27, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003305.

Ashmore:2025:CCC

[310] Rylo Ashmore, Danny Dyer, Trent Marbach, and Rebecca Milley. Cuts, cats, and complete graphs. Theoretical Computer Science, 1050(??): ??, September 27, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003317.

Anselmo:2025:TTA

[311] Marcella Anselmo, Matteo Cavallaro, Maria Madonia, and Carla Selmi. Tiling of toroidal arrays with pictures: Uniqueness, shift-equivalence and undecidability. Theoretical Computer Science, 1050 (??):??, September 27, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003329.

Gao:2025:SHM

[312] Younan Gao and Andrzej Pelc. Sniffing helps to meet: Deterministic rendezvous of anonymous agents in the grid. Theoretical Computer Science, 1050(??):??, September 27, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003330.

Markakis:2025:CWD

[313] Evangelos Markakis and Georgios Papasotiropoulos. On the complexity of winner determination and strategic control in conditional approval voting. Theoretical Computer Science, 1050(??):??, September 27, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003342.

Anonymous:2025:EBbb

[314] Anonymous. Editorial Board. Theoretical Computer Science, 1050(??): ??, September 27, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525004037.

Anonymous:2025:Sd

[315] Anonymous. 27 September 2025. Theoretical Computer Science, 1050(??):??, September 27, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asahiro:2025:MAS

[316] Yuichi Asahiro and Masafumi Yamashita. Minimum algorithm sizes for the gathering and related problems of autonomous mobile robots. Theoretical Computer Science, 1051 (??):??, October 9, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003354.

Aerts:2025:TVB

[317] Diederik Aerts, Roberto Leporini, and Sandro Sozzo. Turing video-based cognitive tests to handle entangled concepts. Theoretical Computer Science, 1051(??):??, October 9, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003366.

Zhang:2025:THC

[318] Yinghao Zhang, Xuanzhe Xiao, and Jianxiong Guo. TransMCS: a hybrid CNN-transformer autoencoder for end-to-end multi-modal medical signals compressive sensing. Theoretical Computer Science, 1051(??): ??, October 9, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003470.

Xia:2025:SRC

[319] Wen Xia, Jorik Jooken, Jan Goedgebeur, and Shenwei Huang. Some results on critical (P_{5}, H)-free graphs. Theoretical Computer Science, 1051 (??):??, October 9, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003494.

Erlebach:2025:PAM

[320] Thomas Erlebach, Nils Morawietz, and Petra Wolf. Parameterized algorithms for multi-label periodic temporal graph realization. Theoretical Computer Science, 1051(??):??, October 9, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003500.

Xia:2025:AAP

[321] Xinlan Xia, Lu Han, and Lili Mei. An approximation algorithm for the parity-constrained k-supplier problem. Theoretical Computer Science, 1051 (??):??, October 9, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003512.

Suzuki:2025:PCW

[322] Takahiro Suzuki, Kei Kimura, Akira Suzuki, Yuma Tamura, and Xiao Zhou. Parameterized complexity of weighted target set selection. Theoretical Computer Science, 1051(??): ??, October 9, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003524.

Aradhya:2025:ROM

[323] Vijeth Aradhya, Seth Gilbert, and Aquinas Hobor. Robust overlays meet blockchains: On handling high churn and catastrophic failures. Theoretical Computer Science, 1051(??): ??, October 9, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003536.

Tao:2025:RMB

[324] Youming Tao, Xiuzhen Cheng, Falko Dressler, Zhipeng Cai, and Dongxiao Yu. Robust matroid bandit optimization: Near-optimal rates under adversarial contamination. Theoretical Computer Science, 1051(??): ??, October 9, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003548.

Dimitrov:2025:SRS

[325] Stoyan Dimitrov. Simple random sampling of binary forests with fixed number of nodes and trees. Theoretical Computer Science, 1051(??): ??, October 9, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003561.

Nouinou:2025:MTC

[326] Hajar Nouinou, Taha Arbaoui, and Alice Yaloui. Minimizing the total completion time for a class of semionline single machine scheduling problems. Theoretical Computer Science, 1051(??):??, October 9, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003603.

Anonymous:2025:EBbc

[327] Anonymous. Editorial Board. Theoretical Computer Science, 1051(??): ??, October 9, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525004104.

Anonymous:2025:Oa

[328] Anonymous. 9 October 2025. Theoretical Computer Science, 1051(??):??, October 9, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goldberg:2025:FIE

[329] Paul W. Goldberg, Kasper Høgh, and Alexandros Hollender. The frontier of intractability for EFX with two agents. Theoretical Computer Science, 1052(??):??, October 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003056.

Li:2025:BBP

[330] Yifei Li, Donghua Yang, and Jianzhong Li. K-balanced biclique partition: Kernelization and efficient algorithms. Theoretical Computer Science, 1052 (??):??, October 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003482.

Wang:2025:AAF

[331] Lu Wang, Qilong Feng, and Jianxin Wang. Approximation algorithms for facility location and k-median with differential privacy. Theoretical Computer Science, 1052(??):??, October 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500355X.

Pshenitsyn:2025:DEP

[332] Tikhon Pshenitsyn. On decidability and expressive power of fusion grammars. Theoretical Computer Science, 1052(??):??, October 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003585.

Bai:2025:QSA

[333] Xujun Bai and Yun Shang. A quantum speedup algorithm for TSP based on quantum dynamic programming with very few qubits. Theoretical Computer Science, 1052(??): ??, October 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003615.

Caroppo:2025:QAO

[334] Susanna Caroppo, Giordano Da Lozzo, and Giuseppe Di Battista. Quantum algorithms for one-sided crossing minimization. Theoretical Computer Science, 1052(??):??, October 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003627.

Gima:2025:CLH

[335] Tatsuya Gima, Tesshu Hanaka, Yasuaki Kobayashi, Yota Otachi, Tomohito Shirai, Akira Suzuki, Yuma Tamura, and Xiao Zhou. On the complexity of list H-packing for sparse graph classes. Theoretical Computer Science, 1052(??):??, October 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003639.

Zhang:2025:CUK

[336] Zhen Zhang, Zhuohang Gao, Limei Liu, Yao Liu, Jie Chen, and Qilong Feng. Clustering under a knapsack constraint: Parameterized approximation for the knapsack median problem. Theoretical Computer Science, 1052(??):??, October 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003640.

Civril:2025:RNC

[337] A. Çivril. Retraction notice to “Corrigendum to ‘A New Approximation Algorithm for the Minimum 2-Edge-Connected Spanning Subgraph Problem’ “ [Theoretical computer science 963 (2023) 113926]. Theoretical Computer Science, 1052(??): ??, October 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003652.

Lucke:2025:MCV

[338] Felicia Lucke. Matching cut and variants on bipartite graphs of bounded radius and diameter. Theoretical Computer Science, 1052(??):??, October 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003676.

Shahar:2025:GCU

[339] Amit Shahar, Daniel Keren, Felipe Gonçalves, and Gal Yehuda. Geometric covering using random fields. Theoretical Computer Science, 1052 (??):??, October 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439752500369X.

Fu:2025:LID

[340] Xinyu Fu, Yitong Yin, and Chaodong Zheng. Locally-iterative (Δ + 1)- coloring in sublinear (in Δ) rounds. Theoretical Computer Science, 1052 (??):??, October 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397525003949.

Adaricheva:2025:CBR

[341] Kira Adaricheva, Lhouari Nourine, and Simon Vilmin. Computing the D-base and D-relation in finite closure systems. Theoretical Computer Science, 1052(??):??, October 19, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003974.

Russyev:2025:FPA

[342] Andriy Russyev. Finiteness problem for automaton groups over a binary alphabet is almost decidable. Theoretical Computer Science, 1052 (??):??, October 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525004165.

Anonymous:2025:EBbd

[343] Anonymous. Editorial Board. Theoretical Computer Science, 1052(??): ??, October 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525004359.

Anonymous:2025:Ob

[344] Anonymous. 19 October 2025. Theoretical Computer Science, 1052(??):??, October 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wojciechowski:2025:URH

[345] P. Wojciechowski and K. Subramani. Unit refutability of horn constraint systems — certification and parallel complexity. Theoretical Computer Science, 1053(??):??, October 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003573.

Chen:2025:SRA

[346] Qun Chen, Qingying Deng, and Kainan Xiang. Subnetwork reliability analysis about complete-transposition graph networks. Theoretical Computer Science, 1053(??):??, October 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397525003597.

Bendotti:2025:ARR

[347] Pascale Bendotti, Luca Brunod-Indrigo, Philippe Chrétienne, and Bruno Escoffier. Approximation results on resource leveling problems. Theoretical Computer Science, 1053(??): ??, October 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397525003688.

Barak-Pelleg:2025:FSM

[348] D. Barak-Pelleg and D. Berend. Fast simulations of the multi-album collector. Theoretical Computer Science, 1053(??):??, October 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397525003706.

Nozaki:2025:REO

[349] Yuta Nozaki, Kunihiro Wasa, and Katsuhisa Yamanaka. Reconfiguration and enumeration of optimal cyclic ladder lotteries. Theoretical Computer Science, 1053(??):??, October 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752500372X.

Amano:2025:DAA

[350] Yuki Amano and Kazuhisa Makino. A 3/4 differential approximation algorithm for traveling salesman problem. Theoretical Computer Science, 1053(??):??, October 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003731.

Busson:2025:CAR

[351] Anthony Busson, Malory Marin, and Rémi Watrigant. Channel allocation revisited through 1-extendability of graphs. Theoretical Computer Science, 1053(??):??, October 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003743.

Adamson:2025:HCT

[352] Duncan Adamson. Harmonious colourings of temporal matchings. Theoretical Computer Science, 1053(??): ??, October 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003755.

Jawhar:2025:BAE

[353] Khaled Jawhar and Evangelos Kranakis. Bike-assisted evacuation of robots on a line with asymmetric S/R communication. Theoretical Computer Science, 1053(??):??, October 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525003767.

Almalki:2025:EGG

[354] Nada Almalki, Siddharth Gupta, and Othon Michail. On the exponential growth of geometric shapes. Theoretical Computer Science, 1053(??): ??, October 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397525003779.

Benso:2025:IPQ

[355] Andrea Benso, Flavio D’Alessandro, and Paolo Papi. On the intersection problem for quantum finite automata. Theoretical Computer Science, 1053(??):??, October 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397525003925.

Daescu:2025:CCA

[356] Ovidiu Daescu and Ka Yaw Teo. Combinatorial characterizations and algorithms for trajectory planning of an articulated robotic probe in three dimensions. Theoretical Computer Science, 1053(??):??, October 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397525003998.

Dubbe:2025:ASP

[357] Thomas Dubbe. The automaticity of the set of primes. Theoretical Computer Science, 1053(??): ??, October 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397525004189.

Xu:2025:MTO

[358] Bo Xu, Zexing Wang, and Zhi-Ping Fan. Model training oriented mobile crowdsourcing tasks: Worker selection and incentive mechanism design. Theoretical Computer Science, 1053(??):??, October 29, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525004499.

Anonymous:2025:EBbe

[359] Anonymous. Editorial Board. Theoretical Computer Science, 1053(??): ??, October 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397525004633.

Anonymous:2025:Oc

[360] Anonymous. 29 October 2025. Theoretical Computer Science, 1053(??):??, October 29, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).